### v1

# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

Abstract
We make three contributions to conformal prediction. First, we propose fuzzy conformal confidence sets that offer a degree of exclusion, generalizing beyond the binary inclusion/exclusion offered by classical confidence sets. We connect fuzzy confidence sets to e-values to show this degree of exclusion is equivalent to an exclusion at different confidence levels, capturing precisely what e-values bring to conformal prediction. We show that a fuzzy confidence set is a predictive distribution with a more appropriate error guarantee. Second, we derive optimal conformal confidence sets by interpreting the minimization of the expected measure of the confidence set as an optimal testing problem against a particular alternative. We use this to characterize exactly in what sense traditional conformal prediction is optimal. Third, we generalize the inheritance of guarantees by subsequent minimax decisions from confidence sets to fuzzy confidence sets. All our results generalize beyond the exchangeable conformal setting to prediction sets for arbitrary models. In particular, we find that any valid test (e-value) for a hypothesis automatically defines a (fuzzy) prediction confidence set.
Keywords: Conformal prediction, prediction sets, e-values, fuzzy confidence sets, certified decisions.
1 Introduction
1.1 Traditional conformal prediction
Suppose we have $n$ exchangeable observations $Z^{n}=(Z_{1},\dots,Z_{n})$. Abstracting away from regressors, the goal of conformal prediction is to use these $n$ observations to construct a confidence set $C_{\alpha}^{Z^{n}}$ that will cover the next observation with a probability of at least $1-\alpha$,
$\Pr(Z_{n+1}\in C_{\alpha}^{Z^{n}})\geq 1-\alpha,$ (1)
assuming $Z^{n+1}=(Z_{1},\dots,Z_{n+1})$ is exchangeable.
To construct such a confidence set, the idea behind conformal prediction is to plug-in hypothetical values $z$ for $Z_{n+1}$ and test whether $(Z_{1},\dots,Z_{n},z)$ is exchangeable at level $\alpha$. Repeating this exercise for each plug-in value $z$ in the sample space, the confidence set is formed by collecting the values $z$ for which the hypothesis is not rejected at level $\alpha$:
$C_{\alpha}^{Z^{n}}=\{z:\textnormal{not reject }(Z_{1},\dots,Z_{n},z)\textnormal{ is exchangeable}\}.$
Figure 1: A traditional and fuzzy conformal confidence set. The solid line represents a traditional level $\alpha=0.05$ confidence set: the points for which this line is zero form the confidence set, and the points at which it equals 11 its complement. The dashed line represents a fuzzy $[0,1]$-valued confidence set, offering a degree of exclusion at each point.
1.2 Fuzzy conformal prediction
A downside of a traditional conformal confidence set is that it is binary in nature: a sample point $z$ either falls inside the confidence set or not. To solve this, we propose fuzzy conformal confidence sets.
For each point $z\in\mathcal{Z}$, a traditional confidence set may be reinterpreted as a function $C_{\alpha}^{Z^{n}}:\mathcal{Z}\to\{0,1\}$, which outputs an inclusion (0) or exclusion (1) for each point $z\in\mathcal{Z}$. A fuzzy confidence set $\widetilde{C}_{\alpha}^{Z^{n}}:\mathcal{Z}\to[0,1]$ generalizes beyond the binary inclusion-exclusion to a degree of exclusion, in the interval $[0,1]$. We illustrate this in Figure 1, where the solid line represents a traditional confidence set and the dashed line a fuzzy confidence set.
For a fuzzy confidence set, the traditional coverage guarantee (1) generalizes to the guarantee that the degree of exclusion of $Z_{n+1}$ is bounded in expectation by $\alpha$:
$\mathrm{E}[\widetilde{C}_{\alpha}^{Z^{n}}(Z_{n+1})]\leq\alpha.$ (2)
For a binary test, this collapses to the classical coverage guarantee (1), since $C_{\alpha}^{Z^{n}}(z_{n+1})=\mathbb{I}\{z_{n+1}\not\in C_{\alpha}^{Z^{n}}\}$.
1.3 E-values and fuzzy confidence
An apparent downside of a fuzzy confidence set is that the ‘degree of exclusion of $z$ at level $\alpha$’ is difficult to interpret. We overcome this by merging the degree of exclusion into the significance level, allowing different points $z$ to be excluded at different significance levels. This relies on the recently established equivalence between fuzzy tests and ‘e-values’ (Koning, 2025b).
To facilitate this, we rescale the confidence set by dividing it by $\alpha$:
$\mathcal{E}^{Z^{n}}(z)=C_{\alpha}^{Z^{n}}(z)/\alpha.$
Due to the rescaling, $\mathcal{E}^{Z^{n}}(z)=1/\alpha$ now corresponds to a full exclusion at level $\alpha$. More generally, we show that the value of a fuzzy confidence set $\mathcal{E}^{Z^{n}}(z)$ at each point $z$ may be directly interpreted as (the reciprocal of) the smallest significance level at which $z$ is excluded, under a generalized Type I error guarantee (Grünwald, 2023a; Koning, 2024). This replaces the ‘degree of exclusion of a point $z$ at level $\alpha$’ by an ‘exclusion at level $\alpha^{\prime}=1/\mathcal{E}^{Z^{n}}(z)$’. Because of the rescaling, the guarantee (2) must be reformulated to
$\mathrm{E}[\mathcal{E}^{Z^{n}}(Z_{n+1})]\leq 1,$
which also enables $0,\infty]-$valued fuzzy confidence sets.
We illustrate fuzzy confidence sets on this evidence scale in Figure 2, which is a rescaled version of Figure 1, now also including a $0,\infty]-$valued fuzzy confidence set. Here, we see that the traditional confidence set only offers an exclusion at level $1/20=0.05$, whereas the fuzzy confidence sets offer exclusions at various significance levels for different points in the sample space. Compared to the traditional confidence set, the unbounded fuzzy confidence set expresses very strong evidence against large values of $z_{n+1}$. The bounded version is capped at evidence corresponding to a rejection at level $1/20=0.05$, but trades this against more evidence for small values of $z_{n+1}$ compared to the unbounded version.
Figure 2: A traditional and fuzzy conformal confidence sets on the evidence scale. The solid line represents a traditional level $\alpha=0.05$ confidence set, the dashed line represents a fuzzy $0,1/\alpha]-$valued confidence set, and the dotted line a fuzzy $0,\infty]-$valued confidence set. An evidence value of $\mathcal{E}^{Z^{n}}(z)=e$ corresponds to an exclusion at significance level $1/e$.
1.4 Optimal prediction sets
Besides generalizing to fuzzy confidence sets, we also contribute to classical non-fuzzy conformal prediction by studying its optimality. In fact, our optimality results generalize beyond exchangeability to prediction confidence sets for arbitrary models.
The key to unlocking our optimality results is to show that a prediction confidence set is equivalent to a single hypothesis test, which is valid for a model $\mathcal{P}$ if and only if the test is valid for the hypothesis $\mathcal{P}$. This means that any test that is valid for a hypothesis $\mathcal{P}$ automatically defines a prediction confidence set.
Next, we use the equivalence between a prediction confidence set and a test to study optimal confidence sets. In particular, we show that a confidence set is of minimal expected size under some measure $\mu$,
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{Z^{n}}[\mu(C_{\alpha})],$
if and only if the corresponding test is most powerful against some alternative hypothesis that incorporates $\mu$. More generally, this measure may be a kernel $\mu^{|Z^{n}}$ that is allowed to depend on $Z^{n}$.
In Section 4, we specialize the optimality to the classical conformal prediction setting where $\mathcal{P}$ is the exchangeable model to derive optimal confidence sets for conformal prediction. We use this to show in what sense classical conformal prediction is optimal, given a choice of conformity score.
We show these results generalize to fuzzy confidence sets, which we show are equivalent to e-values. Optimal fuzzy confidence sets may then be derived from e-values that are optimal against a particular utility function $U$. Classical optimal confidence sets are recovered when using the ‘Neyman–Pearson utility function’ $U:x\mapsto x\wedge 1/\alpha$, which correspond to a Neyman–Pearson most powerful test. This Neyman–Pearson utility function expresses a linear preference up to $1/\alpha$, and no interest in evidence beyond a rejection at the pre-specified level $\alpha$. Other utility functions may be used to express other preferences, which subsequently shape the fuzzy confidence set.
1.5 Certifying subsequent decisions
As a final contribution, treated in Section 5, we explain how a fuzzy confidence set can be used in subsequent decision making. Here, we show how the coverage guarantee offered by a fuzzy confidence set may be inherited by subsequent minimax decisions. This generalizes ideas of Kiyani et al. (2025) from traditional conformal confidence sets to fuzzy confidence sets and general prediction confidence sets.
Most of our findings here build on results of Grünwald (2023b) and Andrews and Chen (2025) to conformal prediction. This relies on the surprising (to us) observation that a fuzzy confidence set is equivalent to (the reciprocal of) an E-posterior; an object recently studied by Grünwald (2023b). This shows that the ‘E-posterior’ is deeply connected to classical hypothesis testing. Furthermore, our perspective unifies the notions of P-certification and E-certification that were recently developed by Andrews and Chen (2025).
1.6 Related literature
To the best of our knowledge, fuzzy confidence sets were first explicitly proposed by Geyer and Meeden (2005) for confidence sets on parameter spaces. Compared to their work, we deliver three innovations. First, Geyer and Meeden (2005) stick to a classical Neyman–Pearson-style power objective, limiting the use of fuzzy confidence sets to discrete data. We generalize to different expected-utility-based power objectives, leading to properly fuzzy confidence sets even in non-discrete data. Second, we show that a ‘degree of inclusion at level $\alpha$’ is equivalent to an exclusion at a different significance level under a generalized Type I error guarantee. Third, we consider fuzzy prediction confidence sets, which are on the sample space instead of on a parameter space.
Vovk (2024) states that precursors of conformal prediction actually relied on e-values (under a different name). He attributes their demise in conformal prediction to the fact that they found obtaining the confidence guarantee more natural with p-values. This is in harmony with our finding that e-values are equivalent to fuzzy confidence sets, showing that e-values play no role in classical (non-fuzzy) confidence sets. Moreover, Vovk (2024) cites the advantage of p-values for constructing conformal predictive distributions. In Remark 2, we show that conformal predictive distributions are (the reciprocal of) defective fuzzy confidence sets. Equivalently, we find that fuzzy conformal confidence sets are (the reciprocal of) conformal predictive distributions that offer an additional desirable joint guarantee, which may be inherited by subsequent decisions.
To the best of our knowledge, the first explicit connection between conformal prediction and e-values is made in the first ArXiv version of Wang and Ramdas (2022) (but omitted in the journal version), where they briefly mention conformal prediction as a potential application of a ‘softrank’ e-value they derive for exchangeability. Koning (2025a) independently derived a more general class of e-values of the form
$\frac{T(Z^{n+1})}{1/(n+1)!\sum_{\pi}T(\pi Z^{n+1})},$ (3)
where $\pi Z^{n+1}$ denotes a permutation $\pi$ of the elements of the tuple $Z^{n+1}$, which nests softrank e-value for a particular choice of the non-negative statistic $T$. Moreover, he treats conformal prediction with e-values in two brief examples, highlighting the potential for post-hoc selection of $\alpha$ (Grünwald, 2023a; Koning, 2024). Soon after, Balinsky and Balinsky (2024) independently discovered (3), using such e-values as a tool to construct traditional conformal confidence sets. More recently, Gauthier et al. (2025) reconnected e-values to the post-hoc selection of $\alpha$, also focusing on e-values of the form (3).
We make several contributions to this line of literature. First, we use the equivalence between e-values and fuzzy testing derived by Koning (2025b) to show that fuzzy confidence sets are precisely what e-values offer to conformal prediction. Second, while such a fuzzy confidence set permits the subsequent post-hoc selection of a single confidence set, we explicitly recommend not to do so, and propose to instead report the entire fuzzy confidence set. Third, and perhaps most important, preceding applications of e-values in conformal prediction rely on e-values of the form (3), but provide little to no guidance on how to select the statistic $T$. This is problematic, as any admissible e-value is of this form (Koning, 2025a). We instead derive optimal e-values for conformal prediction against alternatives that yield ‘small’ fuzzy conformal confidence sets. Finally, the preceding work considers e-values and conformal prediction within the classical exchangeability model, while our work generalizes to prediction confidence sets for arbitrary models.
2 Prediction confidence sets
2.1 Coverage beyond exchangeability
Let $\mathcal{Z}$ denote the sample space of a single observation, and $\mathcal{Z}^{n+1}$ the joint sample space of $(n+1)$ observations. A prediction confidence set is classically defined as a set $C_{\alpha}^{Z^{n}}$ with the level $\alpha>0$ coverage guarantee
$P(Z_{n+1}\in C_{\alpha}^{Z^{n}})\geq 1-\alpha,\textnormal{ for every }P\in\mathcal{P},$ (4)
where $\mathcal{P}$ is some model (collection of distributions) on $\mathcal{Z}^{n+1}$. We say that such a confidence set is valid for $\mathcal{P}$. Conformal prediction is the special case where the model $\mathcal{P}$ is the collection $\mathcal{P}^{\textnormal{exch.}}$ of exchangeable distributions on $\mathcal{Z}^{n+1}$, but we delay specializing to exchangeability in Section 4.
Note that the probability in (4) is over both $Z^{n}$ and $Z_{n+1}$, and so over the entire tuple $Z^{n+1}$. This means that the confidence set may be interpreted as a function of $Z^{n}$. For notational brevity, we write $C_{\alpha}$ to denote this function, and $C_{\alpha}^{z^{n}}$ for the function evaluated at $Z^{n}=z^{n}\in\mathcal{Z}^{n}$.
2.2 The non-coverage event as a test
Our first insight is that a prediction confidence set $C_{\alpha}$ is equivalent to a single test $\phi:\mathcal{Z}^{n+1}\to\{0,1\}$. Moreover, the confidence set is valid for a model $\mathcal{P}$ if and only if the corresponding test is valid for the hypothesis $\mathcal{P}$. A consequence of this result is that any test $\phi$ that is valid for some hypothesis $\mathcal{P}$ automatically produces a prediction confidence set for $\mathcal{P}$.
Due to the importance of this insight, we capture it in Theorem 1. We cover its proof in the main text, as it is simple and constructive.
Theorem 1.
A prediction confidence set $C_{\alpha}$ is equivalent to the test $\phi:\mathcal{Z}^{n+1}\to\{0,1\}$, defined as $\phi(z^{n+1})=\mathbb{I}\{z_{n+1}\in C_{\alpha}^{z^{n}}\}$. The confidence set is valid at level $\alpha$ for $\mathcal{P}$ if and only if the test $\phi$ is.
Proof.
Given a prediction confidence set $C_{\alpha}$, we define the test $\phi$ so that it rejects when $C_{\alpha}^{z^{n}}$ does not cover $z_{n+1}$:
$\phi(z^{n+1})=\mathbb{I}\{z_{n+1}\not\in C_{\alpha}^{z^{n}}\}.$ (5)
Conversely, given the test $\phi$ we may recover the confidence set $C_{\alpha}$ by evaluating it at different points $(z_{1},\dots,z_{n},z_{n+1})$ in the sample space $\mathcal{Z}^{n+1}$:
$C_{\alpha}^{z^{n}}=\{z\in\mathcal{Z}:\phi(z_{1},\dots,z_{n},z)=0\}.$
As we may convert from one to the other without loss, the test $\phi$ and confidence set $C_{\alpha}$ are equivalent.
The equivalence of their validity directly follows from rewriting the coverage guarantee as
$P(Z_{n+1}\not\in C_{\alpha}^{Z^{n}})\leq\alpha,\textnormal{ for every }P\in\mathcal{P},$
[…]
2.3 From confidence width to power
By Theorem 1, constructing a valid prediction confidence set for $\mathcal{P}$ is equivalent to constructing a valid test for $\mathcal{P}$. It then remains to select a test that yields a ‘good’ or even optimal confidence set in some appropriate sense. Here, one may hope that an optimal confidence set corresponds to a test that is most powerful against some alternative. We show this is indeed the case, by deriving the alternative hypothesis under which the test should be most powerful, in order for our confidence set to be ‘as small as possible’.
To formalize what we mean with ‘as small as possible’, let $\mu$ denote some unsigned measure on $\mathcal{Z}$ selected to measure the size of our confidence set. For the sake of exposition, let us assume we have access to the true distribution $P_{*}^{Z^{n}}$ of $Z^{n}$. It then seems natural to formalize our problem as minimizing the expected $\mu$-size of our confidence set under $P_{*}^{Z^{n}}$:
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{P_{*}^{Z^{n}}}[\mu(C_{\alpha}^{Z^{n}})],$
over valid confidence sets $C_{\alpha}$. For example, if $\mathcal{Z}$ is finite then a natural choice for $\mu$ may be the counting measure, which leads to the minimization of the number of elements in our confidence set:
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{P_{*}^{Z^{n}}}[|C_{\alpha}^{Z^{n}}|].$
In Theorem 2, we show that $C_{\alpha}$ minimizes the expected size of the confidence set if and only if its corresponding test $\phi$ is most powerful against the alternative $Q=P_{*}^{Z^{n}}\times\mu$. In the result, we even show that the choice of $\mu$ is allowed to depend on $Z^{n}$.
Theorem 2.
Let $\mu^{|Z^{n}}$ be a probability kernel dependent on $Z^{n}$. Then $C_{\alpha}$ minimizes
$\mathrm{E}^{P_{*}^{Z^{n}}}[\mu^{|Z^{n}}(C_{\alpha}^{Z^{n}})],$
if and only if its corresponding test $\phi$ is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
Proof.
This result follows immediately from the equivalence between the confidence set $C_{\alpha}$ and its test $\phi$:
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{P_{*}^{Z^{n}}}[\mu^{|Z^{n}}(C_{\alpha}^{Z^{n}})]$
$=\operatorname*{arg\,min}_{\phi}\int_{\mathcal{Z}^{n}}\int_{\mathcal{Z}}\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}\,d\mu^{|Z^{n}}\,dP_{*}^{Z^{n}}$
$=\operatorname*{arg\,min}_{\phi}\int_{\mathcal{Z}^{n+1}}(1-\phi)\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{\phi}\int_{\mathcal{Z}^{n+1}}\phi\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{\phi}\mathrm{E}^{Q}[\phi],$
where the third equality follows from the fact that $\mu^{|Z^{n}}$ is a probability kernel. ∎
In practice, the true distribution $P_{*}^{Z^{n}}$ is generally unknown, so that the optimal procedure under the true distribution cannot be used in practice. However, we must stress that the lack of such oracle knowledge does not endanger the validity of the procedure; it only harms its performance. Indeed, if we plug-in some guess for $P_{*}^{Z^{n}}$, then we would still minimize the size of the confidence set in expectation over our guess. As a consequence, we view the choice of plug-in for $P_{*}^{Z^{n}}$ as expressing the distribution of $Z^{n}$ under which we wish the confidence set to be small in expectation.
The fact that we have no oracle knowledge of $P_{*}^{Z^{n}}$ may be viewed as special case of a general problem in hypothesis testing: we can rarely hope to have a test that is optimal uniformly against every distribution in a composite alternative hypothesis. While there is no general-purpose tool to truly ‘solve’ this problem, it is well-studied and solutions exist for special cases. For example, in Section 2.4 we may get around this issue using the Hunt-Stein theorem. In Section 4, we use other techniques to reduce the hypothesis by using the structure of the problem. In case there is no structure available, a general-purpose tool is to construct a valid test against several (or all) possible guesses of $P_{*}^{Z^{n}}$, and choose $\phi$ as some desired probability mixture over these tests.
2.4 Examples: Gaussian location
In this section, we illustrate our methodology in Gaussian models. These examples may be viewed as an illustration, to show that our framework naturally yields familiar prediction confidence sets. In Example 1, $\mathcal{P}$ contains a single distribution and the observations are independent, so that there is nothing to learn from the $n$ observations; we indeed recover a classical confidence set. In Example 2, we find that our optimal confidence set coincides with the familiar prediction set for the Gaussian location family, which is normally derived through finding a pivot and imposing symmetry of the confidence set. In Example 3, we showcase the use of a kernel $\mu^{\mid Z^{n}}$.
Example 1 (i.i.d. simple Gaussian).
In case $\mathcal{P}$ contains just a single distribution, $\mathcal{P}=\{P\}$, then under independence of $Z^{n}$ and $Z_{n+1}$ there is nothing to learn from the first $n$ observations. As a consequence, the optimal (fuzzy) confidence set $C_{\alpha}$ does not depend on the first $n$ observations $z^{n}\in\mathcal{Z}^{n}$ and so is deterministic.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains a single multivariate Gaussian distribution $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$, for known $\mu\in\mathbb{R}$ and $\sigma>0$. A natural goal would be to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our confidence set in expectation. This would come down to choosing the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\lambda$. Unfortunately, such a non-finite measure causes the optimization problem to be poorly defined. Luckily, this may be resolved by replacing it with the proper alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for $\tau>\sigma$. We choose $\tau>\sigma$ so that the likelihood ratio $\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}$ is a monotone transformation of the one with the Lebesgue measure $\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}$.
For the classical Neyman–Pearson objective $\mathrm{E}^{Q}[\phi]$, it is then straightforward to show that the optimal test is the level $\alpha$ Neyman–Pearson optimal test:
$\phi(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})>\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})>\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{|z_{n+1}-\mu|>\sigma c_{1-\alpha/2}\right\},$
where $c_{1-\alpha/2}$ denotes the standard Gaussian’s $(1-\alpha/2)$-quantile, and the replacement by $\lambda$ comes from the invariance of Neyman–Pearson tests to monotone transformations of the likelihood ratio.
Note that this test is independent of $z^{n}$. Hence, we may simply evaluate it at every $z_{n+1}\in\mathcal{Z}$, for arbitrary $z^{n}$, to invert the test in $\mathcal{Z}$ and obtain the confidence set
$\left[\mu-\sigma c_{1-\alpha/2},\mu+\sigma c_{1-\alpha/2}\right],$ (6)
which is simply the standard two-sided Gaussian confidence interval. Here, we indeed have that the probability $Z_{n+1}$ is in the confidence set is $1-\alpha$.
Example 2 (i.i.d. Composite Gaussian).
In case $\mathcal{P}$ is composite, the first $n$ observations are helpful in narrowing down the true distribution $P^{*}\in\mathcal{P}$. As a consequence, the optimal confidence set $C_{\alpha}^{Z^{n}}$ will now depend on the first observations $Z^{n}$.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains i.i.d Gaussians with known variance $\sigma^{2}>0$, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1}):\mu\in\mathbb{R}\}$. We again wish to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our confidence set in expectation over the true distribution. For this purpose, we pick the proper composite alternative $\mathcal{Q}=\{\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2}):\mu\in\mathbb{R}\}$, where $\tau>\sigma$. Moreover, define $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$, where $\bar{Z}_{n}$ is the sample mean of $Z^{n}$.
We now apply the Hunt-Stein theorem. To do this, observe that both $\mathcal{P}$ and $\mathcal{Q}$ are invariant under the shift-map $x\mapsto x+c1_{n+1}$, $c>0$, $x\in\mathbb{R}^{n+1}$, meaning $\textnormal{Law}(X)\in\mathcal{P}\implies\textnormal{Law}(X+c1_{n+1})\in\mathcal{P}$. By the Hunt-Stein theorem, this means that for every valid test there exists a valid invariant test that is at least as powerful. Every invariant test is a function of a maximal invariant statistic. The vector $(A,B)$ is such a maximal invariant, so we may reduce to tests that depend only on $(A,B)$.
Now $A$ has the same distribution for any $\mu$ under both $\mathcal{P}$ and $\mathcal{Q}$, which drops out in the likelihood ratio. Moreover, $B\sim\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)$ under $\mathcal{P}$ and $B\sim\mathcal{N}(0,\tau^{2}+\tau^{2}/n)$ under $\mathcal{Q}$. This reduces the likelihood ratio to
$\textnormal{LR}^{AB}(A,B)=\frac{d\mathcal{N}(0,\tau^{2}+\tau^{2}/n)}{d\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)}(B)=:\textnormal{LR}^{B}(B).$
For the classical Neyman–Pearson power-objective $\mathrm{E}^{Q}[\phi]$, the optimal e-value is the level $\alpha$ Neyman–Pearson optimal test
$\phi(Z^{n+1})$ $=\mathbb{I}\left\{\textnormal{LR}^{B}(B)>\textnormal{LR}^{B}(c_{1-\alpha/2})\right\},$
$=\mathbb{I}\left\{|Z_{n+1}-\bar{Z}_{n}|>c_{1-\alpha/2}\sigma\sqrt{1+1/n}\right\},$
where $c_{1-\alpha/2}$ again denotes the standard Gaussian’s $(1-\alpha/2)$-quantile.
The resulting confidence set for $Z_{n+1}$ is
$[\bar{Z}_{n}-c_{1-\alpha/2}\sigma\sqrt{1+1/n},\bar{Z}_{n}+c_{1-\alpha/2}\sigma\sqrt{1+1/n}].$
Comparing this to (6), we see that the known mean $\mu$ is replaced by an estimator $\bar{Z}_{n}$, and that we have an additional $\sigma^{2}/n$ pop-up in the variance which is due to the estimation of $\bar{Z}_{n}$.
Example 3 (Autoregressive Gaussian).
Now suppose $\mathcal{P}$ contains a single distribution, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\Sigma)\}$, where $\Sigma_{i,j}=\rho^{|i-j|}$, for all pairs $(i,j)$. This corresponds to an AR(1) model with autoregressive parameter $\rho$.
It may now be desirable to choose the kernel $\mu^{|Z^{n}}=\mathcal{N}({\mu_{n},\tau^{2})}$, centered at the conditional mean $\mu_{n}=\mu+\rho(z_{n}-\mu)$ of the final observation $Z^{n+1}$ given $Z^{n}$, where we apologize for the slight clash of notation of the kernel and mean. This implies a joint measure $Q=\mathcal{N}(\mu 1_{n+1},D\Sigma D^{\prime})$, where $\Sigma$ is the same Toeplitz, and $D=\mathrm{diag}(1,\dots,1,\tau)$, with $\tau>1$. The Neyman–Person most powerful test now equals:
$\phi(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(z_{n+1})>\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{\frac{d\lambda}{d\mathcal{N}(\mu_{n},1)}(z_{n+1})>\frac{d\lambda}{d\mathcal{N}(\mu_{n},1)}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{|z_{n+1}-\mu_{n}|>c_{1-\alpha/2}\right\},$
which may again be interpreted as minimizing the Lebesgue measure of the confidence set. The resulting confidence set becomes
$\left[\mu_{n}-\sigma c_{1-\alpha/2},\mu_{n}+\sigma c_{1-\alpha/2}\right],$
which is simply the standard two-sided Gaussian confidence interval for an AR(1) model.
3 Fuzzy confidence sets
3.1 Fuzzy confidence sets
Fuzzy confidence sets are intended to move beyond the binary inclusion-exclusion $Z_{n+1}\not\in C_{\alpha}^{Z^{n}}$ offered by a classical confidence set $C_{\alpha}^{Z^{n}}$. Fuzzy confidence set were originally proposed by Geyer and Meeden (2005) for confidence sets on parameter spaces, to help in expressing confidence in case of discrete data. Our contribution here is to show that they are of general interest, beyond solving discreteness issues.
To describe a fuzzy confidence set, it is helpful to reinterpret the confidence set $C_{\alpha}^{z^{n}}$ as a map
$C_{\alpha}^{z^{n}}:\mathcal{Z}\to\{0,1\},$
where the event $C_{\alpha}^{z^{n}}(z_{n+1})=1$ indicates that $z_{n+1}$ is fully excluded from the confidence set and $C_{\alpha}^{z^{n}}(z_{n+1})=0$ that it is fuzzy included. The coverage guarantee may then be compactly written as
$\mathrm{E}^{P}[C_{\alpha}^{Z^{n}}(Z_{n+1})]\leq\alpha,\textnormal{ for every }P\in\mathcal{P},$
which may be interpreted as guaranteeing that $C_{\alpha}^{Z^{n}}$ excludes $Z_{n+1}$ with probability at most $\alpha$.
A fuzzy confidence set $\widetilde{C}_{\alpha}^{z^{n}}:\mathcal{Z}\to[0,1]$ generalizes the binary inclusion-exclusion to a degree of exclusion, where $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=1$ means a full exclusion, $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=0$ a full inclusion, and $0<\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})<1$ a partial exclusion. We say that a fuzzy confidence set is valid if the average degree of exclusion of the true value $Z_{n+1}$ is bounded by $\alpha$:
$\mathrm{E}^{P}[\widetilde{C}_{\alpha}^{Z^{n}}(Z_{n+1})]\leq\alpha,\textnormal{ for every }P\in\mathcal{P}.$
In the same way that a binary confidence set is equivalent to a single binary test, as expressed in Theorem 1, a fuzzy confidence set is equivalent to a single fuzzy test $\widetilde{\phi}:\mathcal{Z}^{n+1}\to[0,1]$. Indeed, the fuzzy confidence set $\widetilde{C}_{\alpha}^{z^{n}}$ may be recovered by evaluating the fuzzy test $\widetilde{\phi}$ at different $z\in\mathcal{Z}$ in the sample space
$\widetilde{C}_{\alpha}^{z^{n}}(z)=\widetilde{\phi}(z_{1},\dots,z_{n},z),$
and the fuzzy confidence set is valid if this test $\widetilde{\phi}$ is valid. We delay presenting this result to a slightly more general result in Section 3.2.
Remark 1 (Randomization).
If desired, a binary confidence set may be retrieved from a fuzzy confidence set through randomization. In particular, we may independently draw $U\sim\textnormal{Unif}[0,1]$, and construct the (randomized) confidence set
$C_{\alpha}^{Z^{n}}=\mathbb{I}\left\{\widetilde{C}_{\alpha}^{Z^{n}}\geq U\right\},$
which excludes points for which the degree of exclusion exceeds $U$. This randomized confidence set then still satisfies the classical coverage guarantee, where the coverage probability includes the randomization. We generally do not recommend such a randomized procedure, unless a binary confidence set is practically necessary.
3.2 Inclusion at different levels $\alpha$
A downside of fuzzy confidence sets is that the ‘degree of exclusion at level $\alpha$’ is difficult to interpret. We overcome this by using recent innovations in the e-value literature developed in Koning (2025b), to merge the degree of exclusion into the significance level. Instead of a ‘fuzzy exclusion’, we view a fuzzy confidence set as excluding values $z\in\mathcal{Z}$ at different significance levels.
Without loss of generality, the first step towards this interpretation is to divide everything by the significance level $\alpha$, moving from the $[0,1]$-scale to the $0,1/\alpha]-$scale. We refer to this scale as the evidence scale. To avoid ambiguity, we denote the fuzzy confidence set on the evidence scale by $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,1/\alpha]$, defined as
$\mathcal{E}^{z^{n}}(z)=\frac{\widetilde{C}_{\alpha}^{z^{n}}(z)}{\alpha}.$
Due to the division by $\alpha$, it is said to be valid if
$\mathrm{E}^{P}[\mathcal{E}^{Z^{n}}(Z_{n+1})]\leq 1,$ (7)
which also permits the slight generalization where $\alpha=0$, resulting in a map $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,\infty]$.
A consequence of the rescaling is that a full-exclusion of $z$ at significance level $\alpha$ now corresponds to $\mathcal{E}^{z^{n}}(z)=1/\alpha$ for a $0,1/\alpha]-$valued fuzzy confidence set. For example, suppose that we pre-specify $\alpha=0.05$. Then, $\mathcal{E}^{z^{n}}(z)=20$ corresponds to an exclusion of $z$ at level $1/20=0.05$. A natural question is then whether $\mathcal{E}^{z^{n}}(z)=10$, say, may be interpreted as an exclusion of level $\alpha^{\prime}=1/10=0.1$? Remarkably, it turns out that the answer is yes, if we appropriately handle the fact that this makes the significance level data-dependent (Grünwald, 2023a; Koning, 2024). More generally, a fuzzy confidence set $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,\infty]$ may be interpreted as reporting for each value $z\in\mathcal{Z}$ the smallest significance level at which it is excluded from the confidence set.
We now formalize precisely what we mean with this statement. To start, we move to a different representation of our fuzzy confidence set $\mathcal{E}^{Z^{n}}$ as a collection of its $\alpha$-labeled sublevel sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$, by collecting all $z\in\mathcal{Z}$ for which the evidence is less than $1/\alpha$:
$C_{\alpha}^{Z^{n}}=\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\}.$ (8)
An intermediate observation is that these sets $C_{\alpha}^{Z^{n}}$ are each, individually, valid confidence sets:
$\frac{P(Z_{n+1}\not\in C_{\alpha}^{Z^{n}})}{\alpha}$ $=\frac{P(\mathcal{E}^{Z^{n}}(Z_{n+1})\geq 1/\alpha)}{\alpha}$
$\leq\mathrm{E}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1,$ (9)
where the first inequality follows from Markov’s inequality, and the second inequality follows from the validity of the fuzzy confidence set (7).
Markov’s inequality is generally extremely wasteful. We may close the slack introduced in (9) by Markov’s inequality by instead showing that the family of confidence sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$ together satisfies a much stronger joint coverage guarantee for data-dependent levels. This may be viewed as a generalization of (9) to data-dependent significance levels, stating that (9) must hold ‘on average’ over the data-dependent level. The proof mimics that of Theorem 2 in Koning (2024), applied to the exclusion event $\mathbb{I}\{Z_{n+1}\not\in C_{\alpha}\}$. For completeness, we include it in Appendix A.1. The subsequent Lemma 1 follows immediately from (8).
Theorem 3.
The collection of sublevel sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$ of a valid fuzzy confidence set $\mathcal{E}^{Z^{n}}$ satisfies
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P(Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right]\leq 1,\textnormal{ for every }P\in\mathcal{P},$ (10)
for every data-dependent significance level $\widetilde{\alpha}$.
Lemma 1.
The smallest data-dependent level $\widetilde{\alpha}$ for which $z$ is excluded from $C_{\widetilde{\alpha}}$ equals $\widetilde{\alpha}=1/\mathcal{E}^{Z^{n}}(z)$.
Combining Theorem 3 with Lemma 1 shows that we may indeed directly interpret the value of $1/\mathcal{E}^{Z^{n}}(z)$ as the (smallest) level at which $z$ is excluded, and still maintain a generalized Type I error guarantee as in (10).
We end with Theorem 4, which generalizes the equivalence between prediction confidence sets and tests from Theorem 1 to fuzzy prediction confidence sets and e-values. Its proof is omitted, as it mimics that of Theorem 1, plugging in e-values for tests and fuzzy confidence sets for confidence sets.
Theorem 4.
A fuzzy confidence set $\mathcal{E}^{Z^{n}}:\mathcal{Z}\to[0,\infty]$ is equivalent to an e-value $\varepsilon:\mathcal{Z}^{n+1}\to[0,\infty]$, which is valid for $\mathcal{P}$ if and only if the e-value is valid for $\mathcal{P}$.
Remark 2 (Predictive distributions).
Predictive distributions (Vovk et al., 2017) may be interpreted as defective fuzzy confidence sets. Indeed, a standard predictive distribution is equivalent to taking a collection of confidence sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$, which are each merely individually valid at data-independent levels as in (9), and reporting for each $z$ the smallest data-dependent level $\alpha$ at which it is excluded. This means predictive distributions contain a mismatch between their guarantee (for data-independent levels) and what they report (an exclusion at a data-dependent level). A consequence is that if we wish to retain the guarantee in subsequent decisions, then we must pre-specify $\alpha$. In turn, we may only use the corresponding confidence set $C_{\alpha}^{Z^{n}}$ from the collection, which renders the remainder of the confidence distribution redundant.
Fuzzy confidence sets resolve this mismatch, as their validity guarantee (10) is over data-dependent levels. As a consequence, our fuzzy confidence sets may also be interpreted as precisely those kind of predictive distributions that do permit us to look at the entire confidence distribution and still retain a guarantee.
3.3 Optimal fuzzy confidence sets
In this section, we study optimal fuzzy confidence sets, which generalizes the optimality theory for the classical binary confidence sets developed in Section 2.3. The classical binary setting from Section 2.3 is equivalent to choosing the binary e-value $\varepsilon:\mathcal{Z}^{n+1}\to\{0,1/\alpha\}$ that maximizes the objective
$\mathrm{E}^{Q}\left[\varepsilon\right].$ (11)
The naive natural fuzzy generalization would be to choose the $0,1/\alpha]-$valued e-value that maximizes this same objective. This indeed minimizes the expected size of the confidence set, if we were to permit randomization as in Remark 1. Moreover, this corresponds to the classical Neyman–Pearson most powerful randomized test for $\mathcal{P}$ against $Q$, and is the kind of objective considered by Geyer and Meeden (2005) in the context of a classical fuzzy confidence set on a parameter space.
However, to truly unlock the potential of fuzzy confidence sets we consider more general expected-utility objectives (Koning, 2025b). To motivate this, note that the objective (11) over $0,1/\alpha$]-valued e-values is equivalent to maximizing the objective
$\mathrm{E}^{Q}[\varepsilon\wedge 1/\alpha],$
over unbounded e-values $\varepsilon:\mathcal{Z}^{n+1}\to[0,\infty]$, because this objective places no value on evidence beyond $1/\alpha$. This reveals that the objective (11) may be viewed as maximizing the expected utility
$\mathrm{E}^{Q}[U(\varepsilon)],$
under the utility function $U:x\mapsto x\wedge 1/\alpha$. This utility function may be interpreted as expressing the quite rigid preference that we prefer more evidence linearly, up to some threshold $1/\alpha$. That is, we have no interest in seeing exclusions at levels smaller than $\alpha$.
As we do not believe this is the true preference of analysts using prediction confidence sets, we extend to more general utility functions $U:[0,\infty]\to[0,\infty]$. Here, we assume the utility function is non-decreasing and concave, which expresses that we care decreasingly much about obtaining more evidence. In the e-value literature, the most popular utility is log-utility $U:x\mapsto\log(x)$ (Grünwald et al., 2023; Larsson et al., 2024). However, more general utilities have also been studied (Koning, 2025b). One example are power-utilities $U:x\mapsto((x^{h})-1)/h$, $h\neq 0$, where the log-utility appears as the limit as $h\to 1$, or bounded versions of such utilities.
In fact, to obtain properly fuzzy confidence sets, we must move away from Neyman–Pearson-style utility functions $U:x\mapsto x\wedge 1/\alpha$. This is because Neyman–Pearson-style utility function generally produce barely-fuzzy confidence sets, with the fuzziness only appearing in discrete settings. We suspect that Geyer and Meeden (2005) did not realize the potential of fuzzy confidence sets outside of discrete data, as they implicitly stuck to such a Neyman–Pearson-style utility. By extending to more general utilities we extend the scope of fuzzy confidence sets to non-discrete settings.
Maximizing the expected utility may be viewed as maximizing the area under the utility we obtain from excluding each $z$ at a certain level, as measured by $\mu^{|Z^{n}}$ in $Q=P_{*}^{n}\otimes\mu^{|Z^{n}}$:
$\mathrm{E}^{Q}\left[\int_{\mathcal{Z}}U(\mathcal{E}^{Z^{n}}(z))\,d\mu^{|Z^{n}}(z)\right].$
We end with a generalization of Theorem 2 for the fuzzy setting. We omit its proof, as it is analogous to that of Theorem 2.
Theorem 5.
Let $\mu^{|Z^{n}}$ be a $Z^{n}$-dependent probability kernel, and $U:[0,\infty]\to[0,\infty]$. Then $\mathcal{E}$ maximizes
$\mathrm{E}^{P_{*}^{Z^{n}}}\left[\int_{\mathcal{Z}}U(\mathcal{E}^{Z^{n}}(z))\,d\mu^{|Z^{n}}(z)\right],$
if and only if $\varepsilon$ is the $U$-expected-utility optimal e-value against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
3.4 Example: fuzzy Gaussian prediction
Figure 3: Optimal fuzzy confidence sets for Neyman–Pearson-utility (solid), log-utility (dotted) and bounded log-utility (dashed) under the simple Gaussian setting (left) and composite Gaussian setting (right) from Example 1, 2 and 4, for $n=3$, $\mu=0$, $\sigma=1$, $\tau=3.5$, $\bar{Z}_{n}=1.44$. Here, the Neyman–Pearson utility and bounded log-utility are both bounded at $1/0.05=20$.
Example 4 (Fuzzy Gaussian).
In Section 2.4, we implicitly considered optimal fuzzy confidence sets under the classical Neyman–Pearson utility $x\mapsto x\wedge 1/\alpha$. We continue these examples under more general utility functions.
We first consider the log-utility version of Example 1, where we had the simple model $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$ and optimized against the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for known $\mu\in\mathbb{R}$ and $\tau>\sigma>0$. Here, we could choose $\tau\to\infty$ to mimic the Lebesgue measure, as in Section 2.4, but we find that this generally yields undesirable fuzzy confidence sets.
It is well-known that the log-utility-optimal e-value is the likelihood ratio itself. Hence, the log-utility-optimal e-value here is simply the likelihood ratio
$\varepsilon(z^{n+1})=\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})=\mathcal{E}^{z^{n}}(z_{n+1}).$ (12)
For $1/\alpha$-bounded log-utility $U:x\mapsto\log(\varepsilon\wedge 1/\alpha)$, the optimal e-value is the likelihood ratio capped at $1/\alpha$ and ‘boosted’ by some constant $b_{\alpha}$ so that it has expectation exactly 1 (Koning, 2025b):
$\mathcal{E}^{z^{n}}(z_{n+1})=\left(b_{\alpha}\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})\right)\wedge 1/\alpha.$
In the left panel of Figure 3, we display these fuzzy confidence sets next to the traditional confidence set from Example 1.
For the composite setting introduced in Example 2, we had derived the likelihood ratio
$\textnormal{LR}(A,B)=\frac{d\mathcal{N}(0,\tau^{2}+\tau^{2}/n)}{d\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)}(B),$
where $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$. Fixing $\bar{Z}_{n}$ and evaluating this likelihood ratio at plug-in values $z_{n+1}\in\mathbb{R}$ for $Z_{n+1}$, we recover the log-optimal fuzzy confidence set:
$\mathcal{E}^{Z^{n}}(z_{n+1})=\frac{d\mathcal{N}(\bar{Z}_{n},\tau^{2}+\tau^{2}/n)}{d\mathcal{N}(\bar{Z}_{n},\sigma^{2}+\sigma^{2}/n)}(z_{n+1}),$ (13)
which relies on the fact that the Hunt-Stein theorem goes through for log-optimal e-values as proven by Pérez-Ortiz et al. (2022). To the best of our knowledge, this has not been proven yet for general expected-utility-optimal e-values, but we would be surprised if it does not go through in general.
Comparing the fuzzy confidence set in (13) to (12), we see that this likelihood ratio is centered at $\bar{Z}_{n}$ instead of $\mu$ and the variance in both numerator and denominator is inflated by an additional $\tau^{2}/n$- and $\sigma^{2}/n$-term due to the estimation of $\mu$ by $\bar{Z}_{n}$. We illustrate this fuzzy confidence set in the right panel of Figure 3, along its bounded version and the Neyman–Pearson-utility variant from Example 2.
4 Optimal conformal prediction
In this section, we specialize our methodology to conformal prediction, where $\mathcal{P}$ is the class of exchangeable distributions on $\mathcal{Z}^{n+1}$. We cover both optimality for classical conformal confidence sets and more general fuzzy conformal confidence sets.
As we showed in Section 3.2, a fuzzy conformal confidence set is equivalent to an e-value $\varepsilon$, which is valid for $\mathcal{P}$ when $\varepsilon$ is a valid e-value for $\mathcal{P}$. Moreover, in Section 3.3 we show that the fuzzy confidence set is optimal if $\varepsilon$ is optimal for some expected-utility target $\mathrm{E}^{Q}[U(\varepsilon)]$. As a consequence, we may reduce the problem of studying optimal (fuzzy) conformal prediction sets to studying optimal e-values for exchangeability. Optimal e-values for exchangeability (and more general group invariance) have recently been developed in Koning (2025a).
4.1 Background: orbits & exchangeability
Let $\Pi(n+1)$ denote the group of permutations on $n+1$ elements. This group acts on $\mathcal{Z}^{n+1}$ by permuting the elements of each tuple $(z_{1},\dots,z_{n+1})\in\mathcal{Z}^{n+1}$. The group action partitions $\mathcal{Z}^{n+1}$ into a set $\mathcal{O}^{n+1}$ of disjoint orbits $O\in\mathcal{O}^{n+1}$. For a given $z^{n+1}\in\mathcal{Z}^{n+1}$, its orbit is the set of all its permutations:
$O(z^{n+1})=\{z\in\mathcal{Z}^{n+1}:z=\pi z^{n+1},\textnormal{ for some }\pi\in\Pi(n+1)\}.$
On each orbit, we designate a single tuple as its orbit representative, and we define the map $\cdot$] as the map that carries any point on the orbit to this representative.
We say that a random variable $Z^{n+1}$ on $\mathcal{Z}^{n+1}$ is exchangeable if
$Z^{n+1}\overset{d}{=}\pi Z^{n+1},\textnormal{ for every }\pi\in\Pi(n+1).$
Note that this is actually a statement about its distribution $P$, which we say is exchangeable if it is the law of an exchangeable random variable. We use $\mathcal{P}^{\textnormal{exch.}}$ to denote the collection of all exchangeable distributions on $\mathcal{Z}^{n+1}$.
Throughout, we use the following well-known result.
Lemma 2.
$P$ is exchangeable if and only if $P^{|O}=\textnormal{Unif}(O)$, for every $O\in\mathcal{O}^{n+1}$.
4.2 Reduction to hypotheses on orbits
The key idea in Koning (2025a) to obtain an optimal e-value for exchangeability is to partition the problem into a simpler problem on each orbit $O\in\mathcal{O}^{n+1}$. For this purpose, we consider a conditional e-value $\varepsilon_{O}:O\to[0,\infty]$ on each orbit $O\in\mathcal{O}^{n+1}$, defined as the restriction of our e-value $\varepsilon$ to $O$:
$\varepsilon_{O}(z^{n+1})=\varepsilon(z^{n+1}),\textnormal{ for every }z^{n+1}\in O.$
Passing to the collection of these conditional e-values is lossless, as the test $\varepsilon$ may be recovered from the conditional tests, by evaluating the conditional e-value at the data orbit $O(Z^{n+1})$:
$\varepsilon(Z^{n+1})=\varepsilon_{O(Z^{n+1})}(Z^{n+1}).$
The motivation for shattering our e-value into a collection of conditional e-values is Theorem 6, specialized from Theorem 5 in Koning (2025a). It shows that if we construct an optimal conditional e-value $\varepsilon_{O}$ on each orbit $O\in\mathcal{O}^{n+1}$, then the corresponding unconditional e-value $\varepsilon$ is optimal.
Theorem 6 (Koning (2025a)).
If $\varepsilon_{O}$ is valid for $\textnormal{Unif}(O)$ and optimal against $\mathrm{E}^{Q^{|O}}[U(\varepsilon)]$, for every $O$, then $\varepsilon$ is valid for $\mathcal{P}^{exch.}$ and optimal against $\mathrm{E}^{Q}[U(\varepsilon)]$. Moreover, this e-value $\varepsilon$ is optimal uniformly in mixtures over $Q^{|O}$.
4.3 Further reduction of hypotheses
By Theorem 6, it remains to derive an (orbit-conditional) optimal e-value for $\textnormal{Unif}(O)$ against $Q^{|O}$. In this section, we use the available structure of $Q$ to reduce the hypothesis even further.
Recall that we intend to optimize against an alternative of the form $Q=P_{*}^{Z^{n}}\otimes Q^{Z_{n+1}\mid Z^{n}}$, where $P_{*}^{Z^{n}}$ is the ‘true’ law of $Z^{n}$. While we do not know this law of $Z^{n}$, the fact that the true law of $Z^{n+1}$ is exchangeable implies $Z^{n}$ is exchangeable. This allows us to pin $Q^{|O}$ up to the marginal distribution $Q^{Z_{n+1}|O}$ of $Z_{n+1}$, as captured in Proposition 1. Its proof is given in Appendix A.2.
In words, Proposition 1 states that a draw from $Q^{|O}$ is equivalent to first sampling $Z_{n+1}$ using some distribution over the elements of the tuple $Z^{n+1}]$, and subsequently sampling $Z^{n}$ from the remainder uniformly without replacement.
Proposition 1.
If $Z^{n}$ is exchangeable under $Q$, then
$Q^{Z^{n}|O,Z^{n+1}}=\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}),$
where with $\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\})$ we mean uniformly sampled without replacement from the tuple $Z^{n+1}]$ with $Z_{n+1}$ removed.
As a consequence of Proposition 1, the conditional likelihood ratio $\textnormal{LR}^{|O}$ given $O$ reduces to
$\textnormal{LR}^{|O}(z^{n+1})=\frac{dQ^{|O}}{d\textnormal{Unif}(O)}(z^{n+1})$ (14)
$=\frac{d(Q^{Z_{n+1}\mid O}\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}{d(\textnormal{Unif}[Z^{n+1}]\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}(z^{n+1})$ (15)
$=\frac{dQ^{Z_{n+1}\mid O}}{d\textnormal{Unif}[Z^{n+1}]}(z_{n+1})=(n+1)\times Q^{Z_{n+1}\mid O}(z_{n+1})$ (16)
$=:\textnormal{LR}^{Z_{n+1}|O}(z_{n+1}),$ (17)
where $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}:[z^{n+1}]\to\mathbb{R}_{+}$ is a function only of the final element $z_{n+1}$ of the tuple $z^{n+1}$.
To reduce the hypotheses, we may use the finding of Koning (2025b) that expected-utility-optimal e-values are non-decreasing functions of likelihood ratios. As the likelihood ratio reduces to a likelihood ratio of $\textnormal{Unif}[Z^{n+1}]$ to $Q^{Z_{n+1}\mid O}$, this means the optimal problem implicitly reduces to testing the hypotheses
$H_{0}^{O}$ $:Z_{n+1}\sim\textnormal{Unif}[Z^{n+1}],$
$H_{1}^{O}$ $:Z_{n+1}\sim Q^{Z_{n+1}\mid O},$
where $\textnormal{Unif}[Z^{n+1}]$ is well-defined because the orbit representative $Z^{n+1}]$ is uniquely identified by the orbit $O$.
The overall reduction of the alternative from an arbitrary probability measure $Q$ on $\mathcal{Z}^{n+1}$ to $Q^{Z_{n+1}\mid O}$ greatly reduces the complexity of the object we must specify.
Remark 3 (Conformal with orbit-level ranks).
As optimal e-values only depend on the position of $Z_{n+1}$ on $Z^{n+1}]$, we may equivalently express all the results in terms of the rank of $Z_{n+1}$ among $Z^{n+1}]$.
To define a rank for arbitrary data, we may define it relative to the orbit representative $Z^{n+1}]$. In particular, we define $\textnormal{Rank}(Z^{n+1})$ as the index at which each element in the tuple $Z^{n+1}$ needs to be placed to obtain the representative $Z^{n+1}]$ of $O$. This specializes to the conventional rank for data that has a natural order, if $Z^{n+1}]$ is sorted in ascending order.
Let $R_{n+1}^{O}$ denote last element of the tuple $\textnormal{Rank}(Z^{n+1})$, which corresponds to the rank of $Z_{n+1}$ among $Z^{n+1}]$. We may then equivalently express the reduced hypotheses as
$H_{0}^{O}$ $:R_{n+1}^{O}\sim\textnormal{Unif}\{1,\dots,n\},$
$H_{1}^{O}$ $:R_{n+1}^{O}\sim Q_{R}^{O},$
where $Q_{R}^{O}$ is some arbitrary orbit-dependent distribution on $\{1,\dots,n\}$.
In the literature on conformal prediction it is popular to express everything in terms of unconditional ranks (sometimes confusingly named p-values). Using such unconditional ranks exactly discards the information about the orbit.
4.4 Independence under alternative
In Proposition 2 we further specialize from an exchangeable alternative to an alternative with (unconditional) independence. Concretely, the result implies that if we have an alternative of the form $Q=(P_{*}^{Z^{1}})^{n}\times\mu$, then
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})\propto\frac{d\mu}{dP_{*}^{Z^{1}}}(z_{n+1}).$
If we interpret $\mu$ as some reference measure, a consequence of this result is that optimal e-values will be decreasing in the inverse of the ‘true’ $P_{*}^{Z^{1}}$-density of $z_{n+1}$. We believe this makes intuitive sense, as values of $z_{n+1}$ that have low density may be interpreted as being ‘outlying’ or ‘non-conforming’.
Proposition 2.
Suppose that $Q$ is so that $Y^{n+1}$ is independent and $Y^{n}$ is i.i.d., with $Q^{Z_{n+1}}$ and $Q^{Z_{1}}$ mutually absolutely continuous. Then,
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})$
$=\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{n+1})\Bigg{/}\left(\frac{1}{(n+1)}\sum_{i=1}^{n+1}\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{i})\right).$
4.5 Optimal fuzzy conformal
Now that we have reduced the alternative, we may derive expected-utility-optimal e-values for conformal prediction.
Under some regularity conditions on the utility function $U$, we may obtain a clean and interpretable result, which we present in Theorem 7. Note that the conditional likelihood ratio is central in this optimality result. The result follows our reduction to the hypotheses $H_{0}^{O}$ and $H_{1}^{O}$, and a combination of Theorem 6 with Theorem 1 in Koning (2025b) on the shape of expected-utility-optimal e-values.
Theorem 7.
Assume $U$ is concave, non-decreasing, differentiable, $U(0)=0$, with continuous and strictly decreasing derivative $U^{\prime}$, which has limits $U^{\prime}(0)=\infty$ and $U^{\prime}(\infty)=0$. Then, an expected utility-optimal e-value is given by
$\varepsilon^{U}(z^{n+1})=(U^{\prime})^{-1}\left(\lambda^{O(z^{n+1})}\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\right),$
where $\lambda^{O}\geq 0$ is an orbit-dependent normalization constant. This e-value is expected-utility-optimal, uniformly in any mixture over orbits $O$ and any true exchangeable distribution of $Z_{n}$.
Example 5 (Generalized-mean-optimality).
An example of a utility function that has been primarily studied in the e-value literature is $U:x\mapsto\log(x)$, which yields the conditional likelihood ratio itself as the optimal e-value:
$\varepsilon^{\log}(z^{n+1})=\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}).$
This generalizes to power-utility $U:x\mapsto(x^{h}-1)/h$, $h<1$, where the optimizer equals
$\varepsilon^{h}(Y_{n+1})=\frac{(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}{\frac{1}{n+1}\sum_{j=1}^{n+1}(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}.$
The parameter $h$ can be interpreted as tuning how ‘risky’ the e-value is, where large values of $h$ (close to 1) yield a risky e-value and small values ($h\to$-$\infty$) a less risky e-value. Indeed, assuming $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}>0$, then the $h\to$-$\infty$ limit is the no-risk e-value
$\lim_{h\to-\infty}\varepsilon^{h}\equiv 1.$
Moreover, assuming $\textnormal{LR}^{Z_{n+1}|O}$ is uniquely maximized at $z_{*}^{O}$, the $h\to 1$ limit is the most risky e-value that goes all-in on a single element:
$\lim_{h\to 1}\varepsilon^{h}(z^{n+1})=\begin{cases}n+1,&\textnormal{ if }z_{n+1}=z_{*}^{O(z^{n+1})},\\ 0,&\textnormal{ otherwise}.\end{cases}$
Example 6 (Clipping and dampening e-values).
For subsequent decision making, as treated in Section 5, it may be desirable to place a lower bound $0<b\leq 1$ on our e-values, such as $b=0.01$ or $b=0.1$. This prevents settings in which a single zero-valued e-value may come to dominate subsequent decision making due to the minimax nature of the decisions. The log-optimal e-value with such a lower bound is
$\varepsilon(z^{n+1})=\left(\lambda\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\right)\vee b,$
where $\lambda$ is some normalization constant that ensures $\varepsilon$ is an exact e-value. This may be interpreted as ‘clipping’ the likelihood ratio from below, and simultaneously shrinking it by $\lambda$ to ensure it remains a valid e-value. This may be proven analogously to the proof of upper-bounded e-values covered in Koning (2025b).
An alternative strategy to ‘dampen’ e-values is introduced by Grünwald (2023b). There, he proposes to take an e-value $\varepsilon$ and dampen it to the e-value $b+(1-b)\varepsilon$ that may be interpreted as a $b$-mixture of the e-value and the constant 1. If $\varepsilon$ is the log-utility maximizer, then we find $b+(1-b)\varepsilon$ implicitly maximizes the utility function
$U(\varepsilon)=\log([\varepsilon-b]\vee 0).$
4.6 Optimal classical conformal
In Theorem 8, we show how our optimality framework nests classical conformal prediction as a special case if we choose the Neyman–Pearson utility function $U_{\alpha}^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, $\alpha\in(0,1)$. This shows our result nests classical conformal prediction for any non-conformity score that is a monotone function of $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}=(n+1)\times Q^{Z_{n+1}|O(z^{n+1})}$.
Theorem 8.
Write $O^{\prime}$ for $O(z^{n+1})$. A $U_{\alpha}^{\textnormal{NP}}$-optimal e-value is
$\varepsilon^{\textnormal{NP}}(z^{n+1})=\begin{cases}1/\alpha,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})>c_{\alpha}^{O^{\prime}},\\ k^{O^{\prime}},\hskip-11.38092pt&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})=c_{\alpha}^{O^{\prime}},\\ 0,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})<c_{\alpha}^{O^{\prime}},\end{cases}$
where $c_{\alpha}^{O^{\prime}}$ is the $\alpha$ upper-quantile of $\textnormal{LR}^{|O^{\prime}}(z^{n+1})$ under $\textnormal{Unif}[Z^{n+1}]$, and $k^{O^{\prime}}$ is some orbit-dependent constant that ensures $\varepsilon^{\textnormal{NP}}(Z^{n+1})$ is an exact e-value. This e-value is optimal uniformly in any mixture over orbits, and in any monotone transformation of the likelihood ratio.
To make this more interpretable, we provide Corollary 1, which combines Theorem 8 with our result for the i.i.d. setting in Proposition 2 for minimizing the expected confidence set
$\mathrm{E}^{P_{*}^{Z^{n}}}[\mu(C_{\alpha}^{Z^{n}})].$
Using $f=dP_{*}^{Z_{n+1}}/d\mu$ to denote the density of the true distribution, Corollary 1 shows that classical conformal prediction is optimal in this sense when the conformity score $s$ is chosen as a monotone function of the density $f$. Conversely, given a conformity score $s$ and measure $\mu$, classical conformal prediction is optimal uniformly in the class of true distributions for which $s$ is a monotone function of the corresponding density $f$.
Corollary 1.
Suppose $P_{*}^{Z^{n+1}}$ is i.i.d., and let $s:\mathcal{Z}\to\mathbb{R}_{+}$ denote a conformity score. Then, classical conformal prediction minimizes the expected confidence set size for any measure $\mu$ and true distribution $P_{*}^{Z^{n+1}}$ that satisfies
$s=m\left(\frac{dP_{*}^{Z_{n+1}}}{d\mu}\right),$
where $m:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is some monotone function.
5 Conformal certified decisions
In this section, we show how (fuzzy) prediction confidence sets may be used in subsequent decision making.
We consider a decision maker who is faced with selecting a decision $d$ in a decision space $\mathcal{D}$ while facing uncertainty captured by a random variable $Z_{n+1}$ in the outcome space $\mathcal{Z}$. When selecting the decision $d$, they would ideally like to minimize the ‘oracle loss’ $L:\mathcal{D}\times\mathcal{Z}\to\mathbb{R}$ associated with the decision under the true outcome of $Z_{n+1}$: $\min_{d\in\mathcal{D}}L(d,Z_{n+1})$. Unfortunately, the outcome of $Z_{n+1}$ is not yet available at the time of decision making. Instead, the decision maker relies on a (fuzzy) confidence set for $Z_{n+1}$ to aid them in their decision making process.
5.1 Confidence sets and as-if decisions
We begin by presenting the case where the decision maker receives a conventional non-fuzzy confidence set, before extending to fuzzy confidence sets. For these non-fuzzy confidence sets, this was recently developed byAndrews and Chen (2025) and Kiyani et al. (2025).
Suppose the decision maker is given a valid level $(1-\alpha)$ confidence set $C_{\alpha}^{Z^{n}}\subseteq\mathcal{Z}$ for $Z_{n+1}$ under $\mathcal{P}$. We define the as-if decision $\delta(C_{\alpha}^{Z^{n}})$, as the decision that minimizes the worst-case loss over all possible realizations of $Z_{n+1}$ in the confidence set $C_{\alpha}^{Z^{n}}$:
$\delta(C_{\alpha}^{Z^{n}}):=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in C_{\alpha}^{Z^{n}}}L(d,z),$ (18)
where we assume the minimum exists (Manski, 2021). Note that if the confidence set happens to only contain the true future realization of $Z_{n+1}$, then this action would minimize the oracle loss.
Kiyani et al. (2025) and Andrews and Chen (2025) show that this as-if decision $\delta(C_{\alpha}^{Z^{n}})$ enjoys a $(1-\alpha)$-confidence lower bound on the to-be-realized loss:
$P[L(\delta(C_{\alpha}^{Z^{n}}),Z_{n+1})\leq R(\delta(C_{\alpha}^{Z^{n}}))]\geq 1-\alpha,$
for every $P\in\mathcal{P}$, where
$R(C_{\alpha}^{Z^{n}}):=\sup_{z\in C_{\alpha}^{Z^{n}}}L(\delta(C_{\alpha}^{Z^{n}}),z).$
is the worst-case loss associated with this action if $z\in C_{\alpha}^{Z^{n}}$.
5.2 Fuzzy confidence sets
We now move to fuzzy confidence sets, and showcase two distinct ways in which they generalize the as-if decision making framework: as-if decisions under the post-hoc selection of the confidence level $\alpha$, or, more generally, minimax decisions over a loss function weighted by the fuzzy confidence set.
Minimax decisions over an evidence-weighted loss function have previously been for classical confidence set on parameter spaces under the name ‘E-posterior’ and ‘E-certification’ (Grünwald, 2023b; Andrews and Chen, 2025). While post-hoc level selection has been explored in the context of conformal prediction before (Koning, 2025a; Gauthier et al., 2025), the idea of post-hoc as-if decisions is novel, to the best of our knowledge. Moreover, we are also the first to consider such weighted decisions outside of a parameter space.
5.2.1 Post-hoc level as-if decisions
A downside of subsequent decisions with non-fuzzy confidence sets, as presented in Section 5.1, is that the decision maker must pre-specify a single confidence level $\alpha$. In this section, we show that a fuzzy confidence set $\mathcal{E}^{Z^{n}}$ enables a decision maker to choose among as-if decisions at each confidence level $\alpha$, and select the decision with the desired confidence level $\alpha$ in an arbitrary data-dependent fashion. Overall, this still provides a post-hoc $\alpha$ confidence guarantee.
In particular, we first construct the sublevel sets
$C_{\alpha}^{Z^{n}}$ $=\left\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\right\}.$
For each sublevel set, we then construct the as-if decision
$\delta(C_{\alpha}^{Z^{n}})=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in C_{\alpha}^{Z^{n}}}L(d,z).$
This produces a collection $(\delta(C_{\alpha}^{Z^{n}}))_{\alpha>0}$ of possible decisions. A decision maker may then select the desired level of confidence $\widetilde{\alpha}$ in an arbitrary data-dependent fashion. The resulting associated decision then satisfies the the confidence guarantee
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P\left(L(\delta(C_{\widetilde{\alpha}}^{Z^{n}},z)\geq R(\delta(C_{\widetilde{\alpha}}^{Z^{n}}))\mid\widetilde{\alpha}\right)}{\widetilde{\alpha}}\right]\leq 1,$ (19)
for every data-dependent $\widetilde{\alpha}$ and $P\in\mathcal{P}$, mimicking the confidence guarantee treated in Theorem 3.
As a traditional (non-fuzzy) level $\alpha$ confidence set may be seen as a special binary fuzzy confidence set, we may technically also use data-dependent levels with traditional confidence sets. However, we then find that if we select $\widetilde{\alpha}>\alpha$ yields the same confidence set as $\alpha$ but with a weaker guarantee $\widetilde{\alpha}$, and if $\widetilde{\alpha}<\alpha$ then the resulting post-hoc confidence set is the entire sample space. In this sense, non-fuzzy level $\alpha$ confidence sets effectively forces the choice $\widetilde{\alpha}=\alpha$.
5.2.2 Weighted loss
Instead of producing a collection of possible as-if decisions under a post-hoc level guarantee, it is also possible to weigh our loss by the values of the fuzzy confidence set.
In particular, we may consider the decision
$\delta(\mathcal{E}^{Z^{n}}):=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in\mathcal{Z}}L(d,z)/\mathcal{E}^{Z^{n}}(z),$
where we downweight values of $z$ against which we have much evidence, as such values are unlikely to be the realization of $Z_{n+1}$. This puts most emphasis on values of $z$ that may plausibly be the realization of $Z_{n+1}$.
The corresponding risk bound is
$R(\mathcal{E}^{Z^{n}}):=\sup_{z\in\mathcal{Z}}L(\delta(\mathcal{E}^{Z^{n}}),z)/\mathcal{E}^{Z^{n}}(z),$
on which our decision enjoys the confidence bound
$\mathrm{E}^{P}\left[\frac{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})}{R(\mathcal{E}^{Z^{n}})}\right]$
$=\mathrm{E}^{P}\left[\frac{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})}{\sup_{z\in\mathcal{Z}}L(\delta(\mathcal{E}^{Z^{n}}),z)/\mathcal{E}^{Z^{n}}(z)}\right]$
$\leq\mathrm{E}^{P}\left[\frac{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})}{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})/\mathcal{E}^{Z^{n}}(Z_{n+1})}\right]$
$=\mathrm{E}^{P}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1.$
Such weighted loss bounds have been considered for confidence sets on parameter spaces by Grünwald (2023b) and Andrews and Chen (2025), but to the best of our knowledge we are the first to consider them on a sample space.
5.3 Connecting as-if and weighted
While the weighted loss-approach for fuzzy confidence sets in Section 5.2.2 and the as-if decision from in Section 5.1 may seem distinct, we show that non-fuzzy approach may be viewed as a special limiting case of the fuzzy angle. This relationship was not yet established before; Andrews and Chen (2025) describes both approaches as distinct paradigms. The connection is not obvious, as it relies on our observation that the E-posterior coined by Grünwald (2023b) is equivalent to a fuzzy confidence set, and so a generalization of a confidence set.
Let $\gamma\in(0,1)$, and define the following fuzzy confidence set for every $z\in\mathcal{Z}$:
$\mathcal{E}^{Z^{n}}(z):=\gamma\frac{1}{\alpha}\mathbb{I}\{z\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\frac{1}{\alpha}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}.$ (20)
Note that when $\gamma\to 1$, then $\mathcal{E}$ converges to non-fuzzy confidence set. Moreover, this is indeed a fuzzy confidence set, since
$\mathrm{E}^{P}[\mathcal{E}^{Z^{n}}(Z_{n+1})]$
$=\gamma\frac{1}{\alpha}P(Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}})$
$+(1-\gamma)\frac{1}{\alpha}(1-P(Z_{n+1}\not\in\mathcal{C}_{\alpha}^{Z^{n}}))$
$=\frac{1}{\alpha}P(Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}})\leq 1.$
Plugging this into definition of the risk bound for a given decision $d$ yields
$\sup_{z}\frac{L(d,z)}{\mathcal{E}^{Z^{n}}(z)}$
$=\sup_{z}\frac{L(d,z)}{\gamma\frac{1}{\alpha}\mathbb{I}\{z\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\frac{1}{\alpha}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}}$
$=\sup_{z}\frac{\alpha}{\gamma}\mathbb{I}\{z\not\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)+\frac{\alpha}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z).$
Now, assuming $\mathcal{C}_{\alpha}^{Z^{n}}$ is non-empty and $L(d,z)$ is bounded, we have for $\gamma$ sufficiently close to 1 that this equals
$\sup_{z}\frac{\alpha}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)=\frac{\alpha}{1-\gamma}\sup_{z\in\mathcal{C}_{\alpha}^{Z^{n}}}L(d,z),$
which is the risk function we minimize for non-fuzzy confidence sets in (18), up to the constant $\alpha/(1-\gamma)$ that is irrelevant for the optimization. This shows that the non-fuzzy ‘as-if’ decision can indeed be viewed as a limiting case of the fuzzy procedure.
Appendix A Omitted proofs
A.1 Proof of Theorem 3
Proof.
For each $C_{\alpha}$, (8) reveals that the smallest data-dependent level $\widetilde{\alpha}$ choice for which $Z_{n+1}$ is excluded is $1/\mathcal{E}^{Z^{n}}(Z_{n+1})$. This implies,
$\frac{\mathbb{I}\{Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\leq\frac{\mathbb{I}\{Z_{n+1}\not\in C_{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}\}}{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}=\mathcal{E}^{Z^{n}}(Z_{n+1}),$
for every $\widetilde{\alpha}$. As a consequence,
$\mathrm{E}_{\widetilde{\alpha}}^{P}$ $\left[\frac{P(Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right$]
[…]
$=\mathrm{E}^{P}\left[\frac{\mathbb{I}\{Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\right]\leq\mathrm{E}^{P}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1.$
A.2 Proof of Proposition 1
Proof.
If $Z^{n+1}$ is exchangeable, then $Z^{n}$ is exchangeable. Exchangeability of $Z^{n}$ means that the distribution of $Z^{n+1}$ is invariant under permutations in $\Pi(n)$, which is equivalent to uniformity on each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$, as $\Pi(n)$ is a compact group. Each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$ is simply the subset of a $\Pi(n+1)$-orbit $O\in\mathcal{O}^{n+1}$ in which the $(n+1)$th observation is fixed. In fact, each $\Pi(n+1)$-orbit may be partitioned into $\Pi(n)$-orbits as subsets, each differing by the $(n+1)$th element. Hence, conditionally on $O\in\mathcal{O}^{n+1}$, each conditional distribution $Q^{|O}$ may be viewed as first sampling this $(n+1)$th element using some distributions (and with it the $\Pi(n)$-orbit), and subsequently sampling uniformly from the $\Pi(n)$-orbit that contains this element. ∎
A.3 Proof of Proposition 2
Proof.
By assumption, $Q=Q_{1}\times Q_{1}\times\cdots\times Q_{1}\times Q_{n+1}$. Define $Q^{(i)}$ be this product measure, but with $Q_{n+1}$ in the $i$th position, and $Q_{1}$ in the remaining positions. By the structure of $Q$,
$\overline{Q}=\frac{1}{|\Pi(n+1)|}\sum_{\pi\in\Pi(n+1)}\pi Q=\frac{1}{n+1}\sum_{i=1}^{n+1}Q^{(i)}.$
By Proposition 3 in Koning (2025a), the conditional likelihood ratio is the restriction of
$\frac{dQ}{d\overline{Q}}$
to $O$. By definition of $\overline{Q}$, we have, for every event AA,
$\overline{Q}(A)$ $=\frac{1}{n+1}\sum_{i=1}^{n+1}Q^{(i)}(A)$
$=\frac{1}{n+1}\sum_{i=1}^{n+1}\int_{A}\frac{dQ^{(i)}}{dQ}\,dQ$
$=\int_{A}\left(\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{(i)}}{dQ}\right)\,dQ,$
so that
$\frac{d\overline{Q}}{dQ}=\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{(i)}}{dQ}.$
Taking the reciprocal gives
$\frac{dQ}{d\overline{Q}}=\frac{1}{\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{(i)}}{dQ}}.$ (21)
Next, by definition of $Q$ and $Q^{(i)}$, we have
$\frac{dQ^{(i)}}{dQ}(z^{n+1})$ $=\frac{\tfrac{dQ_{n+1}}{dQ_{1}}(z_{i})}{\tfrac{dQ_{n+1}}{dQ_{1}}(z_{n+1})}$
Substituting this into (21) and rewriting yields
$\frac{dQ}{d\overline{Q}}(z^{n+1})=\frac{\tfrac{dQ_{n+1}}{dQ_{1}}(z_{n+1})}{\frac{1}{n+1}\sum_{i=1}^{n+1}\tfrac{dQ_{n+1}}{dQ_{1}}(z_{i})}.$