### v2

# Fuzzy Prediction Sets: Conformal Prediction with E-values

Abstract
We make three contributions to conformal prediction. First, we propose fuzzy conformal prediction sets that offer a degree of exclusion, generalizing beyond the binary inclusion/exclusion offered by classical prediction sets. We connect fuzzy prediction sets to e-values to show this degree of exclusion is equivalent to an exclusion at different confidence levels, capturing precisely what e-values bring to conformal prediction. We show that a fuzzy prediction set is a predictive distribution with an arguably more appropriate error guarantee. Second, we derive optimal conformal prediction sets by interpreting the minimization of the expected measure of a prediction set as an optimal testing problem against a particular alternative. We use this to characterize exactly in what sense traditional conformal prediction is optimal, and show how this may generally be used to construct optimal (fuzzy) prediction sets. Third, we generalize the inheritance of guarantees by subsequent minimax decisions from prediction sets to fuzzy prediction sets. All results generalize beyond the conformal setting to prediction sets for arbitrary models. In particular, we find that constructing a (fuzzy) prediction set for a model is equivalent to constructing a test (e-value) for that model as a hypothesis.
Keywords: Conformal prediction, prediction set, e-value, fuzzy confidence set, certified decision.
1 Introduction
1.1 Traditional conformal prediction
Suppose we have $n$ exchangeable observations $Z^{n}=(Z_{1},\dots,Z_{n})$. Abstracting away from covariates, the goal of conformal prediction is to use these $n$ observations to construct a prediction set $C_{\alpha}^{Z^{n}}$ that will cover the next observation with a probability of at least $1-\alpha$,
$\Pr(Z_{n+1}\in C_{\alpha}^{Z^{n}})\geq 1-\alpha,$ (1)
assuming $Z^{n+1}=(Z_{1},\dots,Z_{n+1})$ is exchangeable.
To construct such a prediction set, the idea behind conformal prediction is to plug-in hypothetical values $z$ for $Z_{n+1}$ and test whether $(Z_{1},\dots,Z_{n},z)$ is exchangeable at level $\alpha$. Repeating this exercise for each plug-in value $z$ in the sample space, the prediction set is formed by collecting the values $z$ for which the hypothesis is not rejected at $\alpha$:
$C_{\alpha}^{Z^{n}}=\{z:\textnormal{not reject }(Z_{1},\dots,Z_{n},z)\textnormal{ is exchangeable}\}.$
Figure 1: A traditional and fuzzy conformal prediction set. The solid line represents a traditional level $\alpha=0.05$ prediction set: the points $z_{n+1}$ at which this line is zero form the prediction set, and the points at which it equals 11 its complement. The dashed line represents a fuzzy $[0,1]$-valued prediction set, offering a degree of exclusion at each point.
1.2 Fuzzy conformal prediction
A downside of a traditional conformal prediction set is that it is binary in nature: a sample point $z$ either falls inside the set or not. To tackle this, we propose fuzzy conformal prediction sets.
For each point $z\in\mathcal{Z}$, a traditional prediction set may be reinterpreted as a function $C_{\alpha}^{Z^{n}}:\mathcal{Z}\to\{0,1\}$, which outputs an inclusion (0) or exclusion (1) for each point $z\in\mathcal{Z}$. A fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}:\mathcal{Z}\to[0,1]$ generalizes beyond the binary inclusion-exclusion to a degree of exclusion, in the interval $[0,1]$. We illustrate this in Figure 1, where the solid line represents a traditional prediction set and the dashed line a fuzzy prediction set.
For a fuzzy prediction set, the traditional coverage guarantee (1) generalizes to the guarantee that the degree of exclusion of $Z_{n+1}$ is bounded in expectation by $\alpha$:
$\mathrm{E}[\widetilde{C}_{\alpha}^{Z^{n}}(Z_{n+1})]\leq\alpha.$ (2)
For a binary test, this collapses to the classical coverage guarantee (1), since $C_{\alpha}^{Z^{n}}(z_{n+1})=\mathbb{I}\{z_{n+1}\not\in C_{\alpha}^{Z^{n}}\}$.
1.3 E-values and fuzzy prediction
An apparent downside of a fuzzy prediction set is that the ‘degree of exclusion of $z$ at level $\alpha$’ is difficult to interpret. We overcome this by merging the degree of exclusion into the significance level, allowing different points $z$ to be excluded at different significance levels. This relies on the recently established equivalence between fuzzy tests and ‘e-values’ (koning2025continuoustestingunifyingtests).
To facilitate this, we rescale the prediction set by dividing it by $\alpha$:
$\mathcal{E}^{Z^{n}}(z)=C_{\alpha}^{Z^{n}}(z)/\alpha.$
Due to the rescaling, $\mathcal{E}^{Z^{n}}(z)=1/\alpha$ now corresponds to a full exclusion at level $\alpha$. More generally, we show that the value of a fuzzy prediction set $\mathcal{E}^{Z^{n}}(z)$ at any point $z$ may be directly interpreted as (the reciprocal of) the smallest significance level at which $z$ is excluded, under a generalized Type I error guarantee (grunwald2022beyond; koning2024post). This replaces the ‘degree of exclusion of a point $z$ at level $\alpha$’ by an ‘exclusion at level $\alpha^{\prime}=1/\mathcal{E}^{Z^{n}}(z)$’. Because of the rescaling, the guarantee (2) must be reformulated to
$\mathrm{E}[\mathcal{E}^{Z^{n}}(Z_{n+1})]\leq 1,$
which also enables $0,\infty]-$valued fuzzy prediction sets.
We illustrate fuzzy prediction sets on this evidence scale in Figure 2, which is a rescaled version of Figure 1 that also includes a $0,\infty]-$valued fuzzy prediction set. Here, we see that the traditional prediction set only offers an exclusion at level $1/20=0.05$, whereas the fuzzy prediction sets offer exclusions at various significance levels for different points in the sample space. Compared to the traditional prediction set, the unbounded fuzzy prediction set expresses very strong evidence against large values of $z_{n+1}$. The bounded version is capped at evidence corresponding to a rejection at level $1/20=0.05$, but trades this against more evidence for small values of $z_{n+1}$ compared to the unbounded version.
We show that fuzzy prediction sets may easily be merged, by averaging their pointwise evidence under arbitrary dependence, or multiplying under independence. Applying this averaging to non-fuzzy prediction sets generally produces fuzzy prediction sets.
Figure 2: One traditional and two fuzzy prediction sets on the evidence scale. The solid line represents a traditional level $\alpha=0.05$ prediction set, the dashed line represents a fuzzy $0,1/\alpha]-$valued prediction set, and the dotted line a fuzzy $0,\infty]-$valued prediction set. An evidence value of $\mathcal{E}^{Z^{n}}(z)=a$ corresponds to an exclusion at significance level $1/a$.
1.4 Optimal (fuzzy) prediction sets
Besides generalizing to fuzzy prediction sets, we also contribute to classical non-fuzzy conformal prediction by studying its optimality. In fact, our optimality results generalize beyond exchangeability to prediction sets for arbitrary models.
The key to unlocking our optimality results is to show that a prediction set for a model $\mathcal{P}$’ is equivalent to a single test for the hypothesis $\mathcal{P}$. We subsequently link optimal prediction sets to optimal testing theory. In particular, we show that a prediction set is of minimal expected size under some measure $\mu$,
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{Z^{n}}[\mu(C_{\alpha})],$
if and only if the corresponding test is most powerful against some alternative hypothesis that incorporates $\mu$.
In Section 5, we specialize the optimality to the classical conformal prediction setting where $\mathcal{P}$ is the exchangeable model to derive optimal prediction sets for conformal prediction. We use this to show in what sense classical conformal prediction is optimal, given a choice of conformity score.
These results generalize to fuzzy prediction sets, which are equivalent to e-values. We connect optimal fuzzy prediction sets to expected-utility optimal e-values koning2025continuoustestingunifyingtests. Classical conformal prediction is recovered as a special case, for the ‘Neyman–Pearson utility function’ $U^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, when the e-value corresponds to the Neyman–Pearson most powerful test at level $\alpha$. Other utility functions may be used to express different preferences regarding the value of different amounts of evidence, which subsequently shapes the fuzzy prediction set.
1.5 Certifying subsequent decisions
In Section 6, we explain how a fuzzy prediction set can be used in subsequent decision making. Here, we show how the coverage guarantee offered by a fuzzy prediction set may be inherited by subsequent minimax decisions. This generalizes ideas of kiyani2025decision from traditional conformal prediction sets to fuzzy conformal prediction sets and general prediction sets.
Most of our findings here build on and extend results of grunwald2023posterior and andrews2025certified to conformal prediction. This relies on the surprising (to us) observation that a fuzzy confidence set is equivalent to (the reciprocal of) an E-posterior; an object recently studied by grunwald2023posterior. This shows that the ‘E-posterior’ is deeply connected to classical hypothesis testing. Furthermore, our perspective unifies the notions of P-certification and E-certification that were recently developed by andrews2025certified.
1.6 Covariates and application
In Section 7, we show how our results extend to the setting of covariates. In the covariate setting, the tuple $Z^{n+1}$ can be decomposed into $Z^{n+1}=(X^{n+1},Y^{n+1})$, and we measure the covariate $X_{n+1}$ alongside $Z^{n}$. The goal is then to obtain a prediction set for $Y_{n+1}$. Although the use of covariates is standard in conformal prediction, we defer their inclusion to this later section as they add considerable notational overhead and are not essential to our main results.
We apply our methodology to a classical character-recognition problem in Section 8. Here, we showcase what fuzzy prediction sets offer over classical prediction sets, and study how different utility functions shape the fuzzy prediction sets.
2 Related literature
To the best of our knowledge, fuzzy confidence sets were first explicitly proposed by geyer2005fuzzy for confidence sets on parameter spaces. Compared to their work, we deliver several innovations. First, while conceptually the same type of object, we stress that our work is on fuzzy prediction sets, which may be viewed as confidence sets on a sample space instead of on a parameter space. Second, geyer2005fuzzy present fuzzy confidence sets as a solution to under-coverage of classical Neyman–Pearson optimal confidence sets for discrete data in small samples. We advocate for a much wider use, beyond discrete data, by using a generalization of the Neyman–Pearson framework from the recent e-value literature. Third, we show that a fuzzy ‘degree of exclusion at level $\alpha$’ is equivalent to an exclusion at a different significance level under a generalized Type I error guarantee. We believe this may make fuzzy prediction/confidence sets more palatable, as this makes them easier to interpret and work with. Finally, we also attempt to show the relevance of fuzzy prediction sets by showing how they may be used in subsequent decision making, further motivating their relevance over non-fuzzy prediction sets.
As the literature on conformal prediction and e-values remains fairly small, we attempt to give a complete overview. According to vovk2024conformal, precursors of conformal prediction (Gammerman1998; Vovk1999) actually relied on e-values (under a different name). He attributes their demise in conformal prediction to the fact that prediction sets were more naturally obtained by using p-values. This is in line with our finding that e-values are equivalent to fuzzy prediction sets, which implies that e-values play no fundamental role in classical (non-fuzzy) prediction sets.
The first explicit connection between modern conformal prediction and e-values appears in the 2020 arXiv version of vovk2024conformal, where he exploits the merging properties of e-values to enable cross-conformal prediction by averaging over data splits. Since then, several other works have used e-values as a pragmatic tool to construct conformal prediction sets. For example, in the context of conformal selection, lee2024boosting, lee2025full, lee2025selection and nair2025diversifying compute e-values for (non-)selection events and plug these into the e-BH procedure to control the false discovery rate. gauthier2025values averages conformal e-values over ambiguous ground truths, where observations may carry multiple valid labels simultaneously. Moreover, gauthier2025values also considers e-values to select a prediction set with a fixed number of elements. Our work differs from this line of work, as our intention is to directly report the e-values as output in the form of a fuzzy prediction set, rather than using them as an intermediate step to obtain classical prediction sets.
In the early version, vovk2024conformal was also the first to propose using conformal e-values of the form
$\frac{T(Z_{n+1})}{\frac{1}{n+1}\sum_{i=1}^{n+1}T(Z_{i})},$ (3)
where $T$ is a non-negative statistic. This form was later independently rediscovered in the first arXiv version of wang2022false, koning2023post and balinsky2024enhancing, and proven to be the only admissible type of e-values for conformal prediction by koning2023post. A key issue of the form (3) is that there is little guidance on how to appropriately select this statistic $T$ and gauthier2025values find the outcomes to be highly sensitive to this choice. We provide guidance by introducing a general expected-utility framework to conformal prediction, which enables users to quantify their preference for different degrees of evidence, and allows them to specify a measure $\mu$ that determines the most important regions of the outcome space. This nests classical conformal prediction as a special case for a particular choice of the utility function, as well as the setting studied by vovk2024conformal for a logarithmic choice of utility function and counting measure $\mu$ in classification problems. In our application, we illustrate that log-utility is not always appropriate in the conformal prediction context. This is not surprising, as log-utility is generally considered in long-run sequential data-collection settings instead of the single-batch setting we consider here.
3 Prediction sets
Let $\mathcal{Z}$ denote the sample space of a single observation, and $\mathcal{Z}^{n+1}$ the joint sample space of $(n+1)$ observations. [^1] This is easily generalized to more general joint sample spaces, but we stick to copies of the same sample space for the sake of exposition. A prediction set is classically defined as a set $C_{\alpha}^{Z^{n}}$ with the level $\alpha>0$ coverage guarantee
$P(Z_{n+1}\in C_{\alpha}^{Z^{n}})\geq 1-\alpha,\textnormal{ for every }P\in\mathcal{P},$ (4)
where $\mathcal{P}$ is some model (collection of distributions) on $\mathcal{Z}^{n+1}$. This means the prediction set $C_{\alpha}^{Z^{n}}$ will cover $Z_{n+1}$ with at least $1-\alpha$ probability. We say that such a prediction set is valid for the model $\mathcal{P}$. Notice that the probability in (4) is over both $Z^{n}$ and $Z_{n+1}$, so that the coverage guarantee is marginal over the entire tuple $Z^{n+1}$.
Conformal prediction is the special case where the model $\mathcal{P}$ is the collection $\mathcal{P}^{\textnormal{exch.}}$ of exchangeable distributions on $\mathcal{Z}^{n+1}$, but we delay specializing to exchangeability in Section 5. Moreover, we postpone the inclusion of covariates to Section 7 as their inclusion considerably clutters the notation and is not relevant for the discussion here.
3.1 Slicing a prediction set for $Z^{n+1}$
Figure 3: Abstract illustration of the sample space $\mathcal{Z}^{n+1}$, where the horizontal dimension represents $\mathcal{Z}^{n}$ and the vertical dimension $\mathcal{Z}_{n+1}$. The prediction set $C_{\alpha}$ for $Z^{n+1}$ is shaded white, and its complement $\mathcal{Z}^{n+1}\setminus C_{\alpha}$ grey. The prediction set $C_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ appears as a $Z^{n}$-dependent random slice of $C_{\alpha}$.
The first key insight is that the construction of a prediction set $C_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ is equivalent to constructing a prediction set $C_{\alpha}$ for $Z^{n+1}$. Indeed, the prediction set $C_{\alpha}^{z^{n}}$ can be interpreted as a ‘slice’ of a prediction set $C_{\alpha}$:
$C_{\alpha}^{z^{n}}=\{z_{n+1}\in\mathcal{Z}:(z^{n},z_{n+1})\in C_{\alpha}\}.$ (5)
We illustrate this in Figure 3, where union of the slices $\{C_{\alpha}^{z^{n}}\}_{z^{n}\in\mathcal{Z}^{n}}$ forms the prediction set $C_{\alpha}$. Here, we stress that the prediction set $C_{\alpha}$ is deterministic: the randomness in $C_{\alpha}^{Z^{n}}$ is purely due to the random realization of the point $Z^{n}$ at which it is ‘sliced’.
The validity of $C_{\alpha}$ as a prediction set for $Z^{n+1}$ under the model $\mathcal{P}$ is equivalent to the validity of $C_{\alpha}^{Z^{n}}$ in the sense of (4). Indeed, by (5) we have $\mathbb{I}\{Z^{n+1}\in C_{\alpha}\}=\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}$, so that
$P(Z^{n+1}\in C_{\alpha})=P(Z_{n+1}\in C_{\alpha}^{Z^{n}}),$
for any $P$. As a consequence, we may focus our attention to constructing prediction sets $C_{\alpha}$ for $Z^{n+1}$.
Remark 1 (Slicing on observables).
These arguments may be extended to slicing based on any part or statistic of $Z^{n+1}$ that is observed. We illustrate this in the covariates setting described in Section 7 where $Z_{n+1}=(X_{n+1},Y_{n+1})$ and we observe $X_{n+1}$ alongside $Z^{n}$, to construct a prediction set for $Y_{n+1}$. There, we still construct a prediction set $C_{\alpha}$ for $Z^{n+1}$ but then slice it at $(Z^{n},X_{n+1})$.
3.2 Prediction as a testing problem
A second insight is that a prediction set $C_{\alpha}$ is equivalent to a hypothesis test. This is a powerful observation, because the construction of hypothesis tests is one of the most deeply studied topics in statistics.
Recall from the introduction that we may view a prediction set $C_{\alpha}$ for $\mathcal{P}$ as a map $C_{\alpha}:\mathcal{Z}^{n+1}\to\{0,1\}$, which returns an inclusion (0) or exclusion (1) for each point $z^{n+1}\in\mathcal{Z}^{n+1}$. We choose to interpret this map $C_{\alpha}$ as a test, where 0 represents a non-rejection and 1 a rejection. Coupling this to Figure 3, the white region corresponds to the non-rejection region of this test and the grey region to its rejection region.
The prediction set $C_{\alpha}$ (and by extension $C_{\alpha}^{Z^{n}}$) is valid for the model $\mathcal{P}$ if and only if it is a valid test for the ‘hypothesis’ $\mathcal{P}$, since
$P(Z_{n+1}\not\in C_{\alpha}^{Z^{n}})=P(Z^{n+1}\not\in C_{\alpha})=\mathrm{E}^{P}[C_{\alpha}],$
for every $P\in\mathcal{P}$. A consequence is that any test that is valid for $\mathcal{P}$ automatically produces a valid prediction set for the model $\mathcal{P}$, and vice-versa.
Given the importance of this observation, we capture it in Theorem 1 for future reference.
Theorem 1.
A prediction set $C_{\alpha}^{Z^{n}}$ is equivalent to the test $C_{\alpha}$,
$C_{\alpha}(z^{n+1}):=\mathbb{I}\{z^{n+1}\not\in C_{\alpha}^{z^{n}}\}.$
The prediction set $C_{\alpha}^{Z^{n}}$ is valid at level $\alpha$ for the model $\mathcal{P}$ if and only if this test $C_{\alpha}$ is valid at level $\alpha$ for the hypothesis $\mathcal{P}$.
Remark 2 (Connection to classical testing).
At first glance the connection to testing may seem superficial, but we argue it is not. Recall that in classical hypothesis testing we hypothesize a model $\mathcal{P}$ (often called “the hypothesis”) and subsequently observe data. As we observe the data, we must reject the hypothesized model if the data and model are incompatible. In the construction of a prediction sets the roles are reversed: we fix a model $\mathcal{P}$ as the ground truth and hypothesize data $Z^{n+1}=z^{n+1}$. If the two are incompatible, we must now reject the hypothesized data. Extending this argument: if a part $Z^{n}=z^{n}$ of $Z^{n+1}$ is also observed alongside the model $\mathcal{P}$, we must reject the hypothesized part: $Z_{n+1}=z_{n+1}$.
3.3 Optimal prediction sets and power
By Theorem 1, constructing a prediction set is equivalent to constructing a test. It then remains to select a test that yields a ‘good’ or even optimal prediction set in some appropriate sense. One may hope that an optimal prediction set corresponds to a test that is most powerful against some alternative. We show this is indeed the case, by deriving the alternative hypothesis under which the test should be most powerful in order for our prediction set to be ‘as small as possible’.
To formalize what we mean with ‘as small as possible’, let $\mu$ denote some unsigned measure on $\mathcal{Z}$, selected to measure the size of $C_{\alpha}^{Z^{n}}$. For the sake of exposition, let us temporarily assume we have access to the true distribution $P_{*}^{Z^{n}}$ of $Z^{n}$. It then seems natural to formalize our problem as minimizing the expected $\mu$-size of $C_{\alpha}^{Z^{n}}$ under $P_{*}^{Z^{n}}$:
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{P_{*}^{Z^{n}}}[\mu(C_{\alpha}^{Z^{n}})],$
over valid prediction sets $C_{\alpha}$. For example, if $\mathcal{Z}$ is finite then a natural choice for $\mu$ may be the counting measure, which leads to the minimization of the number of elements in our prediction set:
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{P_{*}^{Z^{n}}}[|C_{\alpha}^{Z^{n}}|].$
In Theorem 2, we show that $C_{\alpha}$ minimizes the expected size of $C_{\alpha}^{Z^{n}}$ if and only if its test-representation is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu$. In the result, we even show that the choice of $\mu$ may depend on $Z^{n}$.
Theorem 2.
Let $\mu^{|Z^{n}}$ be a probability kernel dependent on $Z^{n}$. Then $C_{\alpha}$ minimizes
$\mathrm{E}^{P_{*}^{Z^{n}}}[\mu^{|Z^{n}}(C_{\alpha}^{Z^{n}})],$
if and only if the test $C_{\alpha}$ is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
Proof.
This result follows immediately from the equivalence between the prediction set $C_{\alpha}$ and its test representation $C_{\alpha}$:
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{P_{*}^{Z^{n}}}[\mu^{|Z^{n}}(C_{\alpha}^{Z^{n}})]$
$=\operatorname*{arg\,min}_{C_{\alpha}}\int_{\mathcal{Z}^{n}}\int_{\mathcal{Z}}\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}\,d\mu^{|Z^{n}}\,dP_{*}^{Z^{n}}$
$=\operatorname*{arg\,min}_{C_{\alpha}}\int_{\mathcal{Z}^{n+1}}(1-C_{\alpha})\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{C_{\alpha}}\int_{\mathcal{Z}^{n+1}}C_{\alpha}\,d(P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}})$
$=\operatorname*{arg\,max}_{C_{\alpha}}\mathrm{E}^{Q}[C_{\alpha}],$
where the third equality follows from the fact that $\mu^{|Z^{n}}$ is a probability kernel. ∎
3.4 Handling unknown $P_{*}^{Z^{n}}$
In practice, the true distribution $P_{*}^{Z^{n}}$ is generally unknown. Fortunately, we can just plug-in an educated guess or estimator $\widehat{P}^{Z^{n}}$ for $P_{*}^{Z^{n}}$, based on a separate dataset. The resulting prediction set will still be valid for $\mathcal{P}$. The only consequence is that it is optimal for the estimator $\widehat{P}^{Z^{n}}$, instead of for the true distribution $P_{*}^{Z^{n}}$. This mirrors the classical problem in hypothesis testing that if we do not ‘know’ the true alternative distribution in case of a composite alternative, then we can rarely hope to construct a test that is most powerful against it, except for extremely rare settings where there exist uniformly most powerful tests.
In case there is no structure available, a general-purpose tool is to construct a valid test against several (or all) possible guesses of $P_{*}^{Z^{n}}$, and choose our test as some (probability)-weighted average over these tests. However, such an average would usually take value in $[0,1]$, so that this is an example of a fuzzy test, which we treat in Section 4. Another strategy is to construct an optimal test against such a weighted average over different resulting choices of $Q$.
3.5 Examples: Gaussian location
In this section, we illustrate our methodology in Gaussian models. Here, we show that our framework naturally yields familiar prediction sets.
In Example 1, $\mathcal{P}$ contains a single distribution and the observations are independent, so that there is nothing to learn from the $n$ observations; we indeed recover a classical prediction set. In Example 2, we find that our optimal prediction set coincides with the familiar prediction set for the Gaussian location family, which is normally derived through finding a pivot and imposing symmetry of the prediction set. In Example 3, we showcase the use of a kernel $\mu^{\mid Z^{n}}$. In Example 4, we show what happens if we naively use the classical two-sided z-test to produce a prediction set, to highlight the importance to carefully select the alternative $Q$.
Example 1 (i.i.d. simple Gaussian).
In case $\mathcal{P}$ contains just a single distribution, $\mathcal{P}=\{P\}$, then under independence of $Z^{n}$ and $Z_{n+1}$ there is nothing to learn from the first $n$ observations. As a consequence, the optimal (fuzzy) prediction set $C_{\alpha}$ does not depend on the first $n$ observations $z^{n}\in\mathcal{Z}^{n}$ and so is deterministic.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains a single multivariate Gaussian distribution $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$, for known $\mu\in\mathbb{R}$ and $\sigma>0$. A natural goal would be to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our prediction set in expectation. This would come down to choosing the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\lambda$. Unfortunately, such a non-finite measure causes the optimization problem to be poorly defined. Luckily, this may be resolved by replacing it with the proper alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for $\tau>\sigma$. We choose $\tau>\sigma$ so that the likelihood ratio $\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}$ is a monotone transformation of the one with the Lebesgue measure $\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}$.
For the classical Neyman–Pearson objective $\mathrm{E}^{Q}[C_{\alpha}]$, it is then straightforward to show that the optimal test is the level $\alpha$ Neyman–Pearson optimal test:
$C_{\alpha}(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})>\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})>\frac{d\lambda}{d\mathcal{N}(\mu,\sigma^{2})}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{|z_{n+1}-\mu|>\sigma c_{1-\alpha/2}\right\},$
where $c_{1-\alpha/2}$ denotes the standard Gaussian’s $(1-\alpha/2)$-quantile, and the replacement by $\lambda$ comes from the invariance of Neyman–Pearson tests to monotone transformations of the likelihood ratio.
Note that this test is independent of $z^{n}$. Hence, we may simply evaluate it at every $z_{n+1}\in\mathcal{Z}$, for arbitrary $z^{n}$, to invert the test in $\mathcal{Z}$ and obtain the prediction set
$\left[\mu-\sigma c_{1-\alpha/2},\mu+\sigma c_{1-\alpha/2}\right],$ (6)
which is simply the standard two-sided Gaussian prediction interval. Here, we indeed have that the probability $Z_{n+1}$ is in the prediction set is $1-\alpha$.
Example 2 (i.i.d. Composite Gaussian).
In case $\mathcal{P}$ is composite, the first $n$ observations are helpful in narrowing down the true distribution $P_{*}^{Z^{n+1}}\in\mathcal{P}$. As a consequence, the optimal prediction set $C_{\alpha}^{Z^{n}}$ will now depend on the first observations $Z^{n}$.
To illustrate this, let $\mathcal{Z}=\mathbb{R}$, and suppose our class of distributions contains i.i.d Gaussians with known variance $\sigma^{2}>0$, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1}):\mu\in\mathbb{R}\}$. We again wish to minimize the Lebesgue measure $\lambda(C_{\alpha}^{Z^{n}})$ of our prediction set in expectation over the true distribution. For this purpose, we pick the proper composite alternative $\mathcal{Q}=\{\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2}):\mu\in\mathbb{R}\}$, where $\tau>\sigma$. Moreover, define $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$, where $\bar{Z}_{n}$ is the sample mean of $Z^{n}$.
We now apply the Hunt-Stein theorem. To do this, observe that both $\mathcal{P}$ and $\mathcal{Q}$ are invariant under the shift-map $x\mapsto x+c1_{n+1}$, $c>0$, $x\in\mathbb{R}^{n+1}$, meaning $\textnormal{Law}(X)\in\mathcal{P}\implies\textnormal{Law}(X+c1_{n+1})\in\mathcal{P}$. By the Hunt-Stein theorem, this means that for every valid test there exists a valid invariant test that is at least as powerful. Every invariant test is a function of a maximal invariant statistic. The vector $(A,B)$ is such a maximal invariant, so we may reduce to tests that depend only on $(A,B)$.
Now, $A$ has the same distribution for any $\mu$ under both $\mathcal{P}$ and $\mathcal{Q}$, which drops out in the likelihood ratio. Moreover, $B\sim\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)$ under $\mathcal{P}$ and $B\sim\mathcal{N}(0,\tau^{2}+\tau^{2}/n)$ under $\mathcal{Q}$. This reduces the likelihood ratio to
$\textnormal{LR}^{AB}(A,B)=\frac{d\mathcal{N}(0,\tau^{2}+\tau^{2}/n)}{d\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)}(B)=:\textnormal{LR}^{B}(B).$
For the classical Neyman–Pearson power-objective $\mathrm{E}^{Q}[C_{\alpha}]$, the optimal e-value is the level $\alpha$ Neyman–Pearson optimal test
$C_{\alpha}(Z^{n+1})$ $=\mathbb{I}\left\{\textnormal{LR}^{B}(B)>\textnormal{LR}^{B}(c_{1-\alpha/2})\right\},$
$=\mathbb{I}\left\{|Z_{n+1}-\bar{Z}_{n}|>c_{1-\alpha/2}\sigma\sqrt{1+1/n}\right\},$
where $c_{1-\alpha/2}$ again denotes the standard Gaussian’s $(1-\alpha/2)$-quantile.
The resulting prediction set for $Z_{n+1}$ is
$[\bar{Z}_{n}-c_{1-\alpha/2}\sigma\sqrt{1+1/n},\bar{Z}_{n}+c_{1-\alpha/2}\sigma\sqrt{1+1/n}].$
Comparing this to (6), we see that the known mean $\mu$ is replaced by an estimator $\bar{Z}_{n}$, and that we have an additional $\sigma^{2}/n$ appear in the variance which is due to the uncertainty introduced by the estimation of $\mu$ by $\bar{Z}_{n}$.
Example 3 (Autoregressive Gaussian).
Now suppose $\mathcal{P}$ contains a single distribution, $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\Sigma)\}$, where $\Sigma_{i,j}=\rho^{|i-j|}$, for all pairs $(i,j)$. This corresponds to an AR(1) model with autoregressive parameter $\rho\in(-1,1)$.
It may now be desirable to choose the kernel $\mu^{|Z^{n}}=\mathcal{N}({\mu_{n},\tau^{2})}$, centered at the conditional mean $\mu_{n}=\mu+\rho(z_{n}-\mu)$ of the final observation $Z^{n+1}$ given $Z^{n}$, where we apologize for the slight clash of notation of the kernel and mean. This implies a joint measure $Q=\mathcal{N}(\mu 1_{n+1},D\Sigma D^{\prime})$, where $\Sigma$ is the same Toeplitz, and $D=\mathrm{diag}(1,\dots,1,\tau)$, with $\tau>1$. The Neyman–Person most powerful test now equals:
$C_{\alpha}(z^{n+1})$ $=\mathbb{I}\left\{\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(z_{n+1})>\frac{d\mathcal{N}(\mu_{n},\tau^{2})}{d\mathcal{N}(\mu_{n},1)}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{\frac{d\lambda}{d\mathcal{N}(\mu_{n},1)}(z_{n+1})>\frac{d\lambda}{d\mathcal{N}(\mu_{n},1)}(c_{1-\alpha/2})\right\}$
$=\mathbb{I}\left\{|z_{n+1}-\mu_{n}|>c_{1-\alpha/2}\right\},$
which may again be interpreted as minimizing the Lebesgue measure of the prediction set. The resulting prediction set becomes
$\left[\mu_{n}-\sigma c_{1-\alpha/2},\mu_{n}+\sigma c_{1-\alpha/2}\right],$
which is simply the standard two-sided Gaussian prediction interval for an AR(1) model.
Example 4 (Prediction set based on two-sided z-test).
As highlighted in Theorem 1, any valid test produces a valid prediction set. The most popular test for Gaussian data is perhaps the two-sided z-test, so one may wonder what kind of prediction set it produces. Here, our model is $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$, as this is the null hypothesis for which the two-sided z-test for the mean $\mu$ is valid.
Recall that the two-sided z-test rejects a hypothesis at level $\alpha$ if $\bar{Z}_{n+1}>\mu+\sigma c_{1-\alpha/2}/\sqrt{n+1}$, where $c_{1-\alpha/2}$ is the standard Gaussian’s $(1-\alpha/2)$-quantile. From this rejection event, we then recover the prediction set as the values of $Z_{n+1}=z_{n+1}$ for which the test does not reject:
$[\mu+n(\mu-\bar{Z}_{n})$- $\sqrt{n+1}\sigma c_{1-\alpha/2},$
$\mu+n(\mu-\bar{Z}_{n})+\sqrt{n+1}\sigma c_{1-\alpha/2}].$
As the model $\mathcal{P}$ is the same as the one we considered in Example 1, we may directly compare the prediction sets. In this comparison, we see that the prediction set here depends on $Z^{n}$ and is centered around $\mu+n(\mu-\bar{Z}_{n})$ instead of $\mu$. Moreover, to compensate for the peculiar center, its radius is the much wider $\sqrt{n+1}\sigma c_{1-\alpha/2}$ instead of $\sigma c_{1-\alpha/2}$.
To conclude, while this is technically a perfectly valid prediction set, it is perhaps not the prediction set that one would desire to use in most applications. This highlights the importance of selecting an appropriate $Q$ to express our desire about the prediction set.
4 Fuzzy prediction sets
4.1 From classical to fuzzy prediction sets
Fuzzy prediction sets are intended to move beyond the binary inclusion-exclusion $Z_{n+1}\not\in C_{\alpha}^{Z^{n}}$ offered by a classical prediction set $C_{\alpha}^{Z^{n}}$.
To define a fuzzy prediction set, it is helpful to again consider the test-representation of a prediction set
$C_{\alpha}^{z^{n}}:\mathcal{Z}\to\{0,1\},$
where the event $C_{\alpha}^{z^{n}}(z_{n+1})=1$ indicates that $z_{n+1}$ is excluded from the prediction set and $C_{\alpha}^{z^{n}}(z_{n+1})=0$ that it is included. The coverage guarantee may then be compactly written as
$\mathrm{E}^{P}[C_{\alpha}^{Z^{n}}(Z_{n+1})]\leq\alpha,\textnormal{ for every }P\in\mathcal{P},$
which may be interpreted as guaranteeing that $C_{\alpha}^{Z^{n}}$ excludes $Z_{n+1}$ with probability at most $\alpha$.
A fuzzy prediction set $\widetilde{C}_{\alpha}^{z^{n}}:\mathcal{Z}\to[0,1]$ generalizes the binary inclusion-exclusion to a degree of exclusion, where $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=1$ means a full exclusion, $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=0$ a full inclusion, and $0<\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})<1$ a partial exclusion. We say that a fuzzy prediction set is valid if the average degree of exclusion of the true value $Z_{n+1}$ is bounded by $\alpha$:
$\mathrm{E}^{P}[\widetilde{C}_{\alpha}^{Z^{n}}(Z_{n+1})]\leq\alpha,\textnormal{ for every }P\in\mathcal{P}.$
As captured in Theorem 1 for binary prediction sets, a fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ is a slice of a fuzzy prediction set $\widetilde{C}_{\alpha}:\mathcal{Z}^{n+1}\to[0,1]$ for $Z^{n+1}$:
$\widetilde{C}_{\alpha}^{z^{n}}(z)=\widetilde{C}_{\alpha}(z_{1},\dots,z_{n},z).$
Moreover, the fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ is valid if $\widetilde{C}_{\alpha}$ is valid. We delay presenting this result to Section 4.2, where we present it in a slightly more general form in Theorem 4.
Remark 3 (Randomization).
If desired, a binary prediction set may be retrieved from a fuzzy prediction set through randomization. In particular, we may independently draw $U\sim\textnormal{Unif}[0,1]$, and construct the (randomized) prediction set
$C_{\alpha}^{Z^{n}}=\mathbb{I}\left\{\widetilde{C}_{\alpha}^{Z^{n}}\geq U\right\},$
which excludes points for which the degree of exclusion exceeds $U$. This randomized prediction set then still satisfies the classical coverage guarantee, where the coverage probability includes the randomization. We generally do not recommend such a randomized procedure unless a binary prediction set is practically necessary.
4.2 Inclusion at different levels $\alpha$
A downside of fuzzy prediction sets is that the ‘degree of exclusion at level $\alpha$’ is difficult to interpret. We overcome this by using recent innovations in the e-value literature developed in koning2025continuoustestingunifyingtests, to merge the degree of exclusion into the significance level. Instead of a ‘fuzzy exclusion’, we view a fuzzy prediction set as excluding values $z\in\mathcal{Z}$ at different significance levels.
Without loss of generality, the first step towards this interpretation is to divide everything by the significance level $\alpha$, moving from the $[0,1]$-scale to the $0,1/\alpha]-$scale. We refer to this scale as the evidence scale. To avoid ambiguity, we denote the fuzzy prediction set on the evidence scale by $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,1/\alpha]$, defined as
$\mathcal{E}^{z^{n}}(z)=\frac{\widetilde{C}_{\alpha}^{z^{n}}(z)}{\alpha}.$
Due to the division by $\alpha$, it is now said to be valid if
$\mathrm{E}^{P}[\mathcal{E}^{Z^{n}}(Z_{n+1})]\leq 1,$ (7)
which also permits the slight generalization where $\alpha=0$, resulting in a map $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,\infty]$.
A consequence of this rescaling is that a full-exclusion of $z$ at significance level $\alpha$ now corresponds to $\mathcal{E}^{z^{n}}(z)=1/\alpha$ for a $0,1/\alpha]-$valued fuzzy prediction set. For example, suppose that we pre-specify $\alpha=0.05$. Then, $\mathcal{E}^{z^{n}}(z)=20$ corresponds to an exclusion of $z$ at level $1/20=0.05$. A natural question to ask is whether $\mathcal{E}^{z^{n}}(z)=10$, say, may be interpreted as an exclusion of level $\alpha^{\prime}=1/10=0.1$? Remarkably, it turns out that the answer is yes, if we appropriately handle the fact that this makes the significance level $\alpha^{\prime}$ data-dependent (grunwald2022beyond; koning2024post). More generally, a fuzzy prediction set $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,\infty]$ may be interpreted as reporting for each value $z\in\mathcal{Z}$ the smallest significance level at which it is excluded from the prediction set.
We now formalize precisely what we mean with this statement. To start, we move to a different representation of our fuzzy prediction set $\mathcal{E}^{Z^{n}}$ as a collection of its $\alpha$-labeled sublevel sets $(\overline{C}_{\alpha}^{Z^{n}})_{\alpha>0}$, by collecting all $z\in\mathcal{Z}$ for which the evidence is less than $1/\alpha$:
$\overline{C}^{Z^{n}}=\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\}.$ (8)
An intermediate observation is that these sets $\overline{C}_{\alpha}^{Z^{n}}$ are each, individually, valid prediction sets:
$\frac{P(Z_{n+1}\not\in\overline{C}_{\alpha}^{Z^{n}})}{\alpha}$ $=\frac{P(\mathcal{E}^{Z^{n}}(Z_{n+1})\geq 1/\alpha)}{\alpha}$
$\leq\mathrm{E}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1,$ (9)
where the first inequality follows from Markov’s inequality, and the second inequality follows from the validity of the fuzzy prediction set (7).
Markov’s inequality generally has a lot of slack. In Theorem 3, we close this slack by showing that the family of prediction sets $(\overline{C}_{\alpha}^{Z^{n}})_{\alpha>0}$ together satisfies a much stronger joint coverage guarantee for data-dependent levels. This may be viewed as a generalization of (9) to data-dependent significance levels, stating that (9) must hold ‘on average’ over the data-dependent level.
The proof immediately follows from Theorem 2 in koning2024post, applied to the test $\overline{C}_{\alpha}=\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{\alpha}^{Z^{n}}\}$. For completeness, we include it in Appendix A.1. The subsequent Lemma 1 follows immediately from (8).
Theorem 3.
The collection of sublevel sets $(\overline{C}_{\alpha}^{Z^{n}})_{\alpha>0}$ of a valid fuzzy prediction set $\mathcal{E}^{Z^{n}}$ satisfies
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P(Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right]\leq 1,\textnormal{ for every }P\in\mathcal{P},$ (10)
for every data-dependent significance level $\widetilde{\alpha}$.
Lemma 1.
The smallest data-dependent level $\widetilde{\alpha}$ for which $z$ is excluded from $\overline{C}_{\widetilde{\alpha}}$ equals $\widetilde{\alpha}=1/\mathcal{E}^{Z^{n}}(z)$.
Combining Theorem 3 with Lemma 1 shows that we may indeed directly interpret the value of $1/\mathcal{E}^{Z^{n}}(z)$ as the (smallest) level at which $z$ is excluded, and still maintain a generalized Type I error guarantee as in (10).
We end with Theorem 4, which generalizes Theorem 1 to fuzzy prediction sets on the evidence $0,\infty]-$scale. This (slightly) generalizes the discussion in Section 4.1, as the discussion there technically does not cover fuzzy prediction sets on an unbounded evidence scale $0,\infty]$.
A fuzzy test on the evidence $0,\infty]-$scale is also known as an e-value (koning2025continuoustestingunifyingtests). Hence, in order to construct a fuzzy prediction set $\mathcal{E}^{Z^{n}}$, it suffices to construct an e-value and then slice it at $Z^{n}$.
Theorem 4.
A fuzzy prediction set $\mathcal{E}^{Z^{n}}:\mathcal{Z}\to[0,\infty]$ for $Z_{n+1}$ is a slice
$\mathcal{E}^{z^{n}}(z)=\mathcal{E}(z^{n},z)$
of a fuzzy prediction set $\mathcal{E}$ for $Z^{n+1}$. Such a fuzzy prediction set $\mathcal{E}$ is an e-value, and it is valid for $\mathcal{P}$ if and only if it is a valid e-value for $\mathcal{P}$.
Remark 4 (Predictive distributions).
Predictive distributions (vovk2017nonparametric) may be interpreted as fuzzy prediction sets with weaker error guarantees. Indeed, a standard predictive distribution is equivalent to taking a collection of prediction sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$, which are each merely individually valid at data-independent levels as in (9), and reporting for each $z$ the smallest data-dependent level $\alpha$ at which it is excluded. This means predictive distributions contain a mismatch between their guarantee (for data-independent levels) and what they report (an exclusion at a data-dependent level). A consequence is that if we wish to retain the guarantee then we must pre-specify $\alpha$. In turn, we may only use the corresponding prediction set $C_{\alpha}^{Z^{n}}$ from the collection, which renders the remainder of the predictive distribution redundant.
Fuzzy prediction sets resolve this mismatch, as their validity guarantee (10) is over data-dependent levels. As a consequence, our fuzzy prediction sets may also be interpreted as precisely those kind of predictive distributions that do permit us to look at the entire predictive distribution and still retain a guarantee.
4.3 Optimal fuzzy prediction sets
In this section, we study optimal fuzzy prediction sets. This generalizes the optimality theory for the classical binary prediction sets developed in Section 3.3. Instead of viewing this problem as finding an optimal test, we now consider the more general problem of finding an optimal e-value. The classical binary setting from Section 3.3 is equivalent to choosing the binary valid e-value $\mathcal{E}:\mathcal{Z}^{n+1}\to\{0,1/\alpha\}$ that maximizes the objective
$\mathrm{E}^{Q}\left[\mathcal{E}\right].$ (11)
The naive natural fuzzy generalization would be to choose the $0,1/\alpha]-$valued e-value that maximizes this same objective. This indeed minimizes the expected size of the prediction set, if we were to permit randomization as in Remark 3. Moreover, this corresponds to the classical Neyman–Pearson most powerful randomized test for $\mathcal{P}$ against $Q$, and is the kind of objective considered by geyer2005fuzzy in the context of a classical fuzzy prediction set on a parameter space.
However, to truly unlock the potential of fuzzy prediction sets we consider more general expected-utility objectives (koning2025continuoustestingunifyingtests). To motivate this, note that an optimizer of (11) among $0,1/\alpha$]-valued e-values is also an optimizer of the objective
$\mathrm{E}^{Q}[\mathcal{E}\wedge 1/\alpha],$
over unbounded e-values $\mathcal{E}:\mathcal{Z}^{n+1}\to[0,\infty]$, because this objective places no value on evidence beyond $1/\alpha$. This reveals that the objective (11) may be viewed as maximizing the expected utility
$\mathrm{E}^{Q}[U^{\textnormal{NP}}(\mathcal{E})],$
under the utility function $U^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$. We name this the “Neyman–Pearson utility function”, as classical Neyman–Pearson hypothesis testing can be viewed as collecting evidence under this utility function, which expresses linear valuation of evidence up to some threshold $1/\alpha$. This threshold expresses that we put no additional value on seeing exclusions at levels smaller than $\alpha$.
As we do not believe this Neyman–Pearson utility function expresses the true preference of analysts using prediction sets, we extend to more general utility functions $U:[0,\infty]\to[-\infty,\infty]$. Here, we assume the utility function is non-decreasing and concave, which expresses that we care decreasingly much about obtaining more evidence. See Remark 5 for a discussion on the choice of utility.
To obtain p8roperly fuzzy prediction sets, we must in fact move away from Neyman–Pearson-style utility functions $U:x\mapsto x\wedge 1/\alpha$. This is because a Neyman–Pearson-style utility function generally produces barely-fuzzy prediction sets, with the fuzziness only appearing in discrete settings. We suspect that geyer2005fuzzy did not appreciate the potential of fuzzy prediction sets outside of discrete data, as they implicitly stuck to such a Neyman–Pearson-style utility. By extending to more general utilities we extend the scope of fuzzy prediction sets to non-discrete settings.
Maximizing the expected utility may be viewed as maximizing the area under the utility we obtain from excluding each $z$ at a certain level, as measured by $\mu^{|Z^{n}}$ in $Q=P_{*}^{n}\otimes\mu^{|Z^{n}}$:
$\mathrm{E}^{Q}\left[\int_{\mathcal{Z}}U(\mathcal{E}^{Z^{n}}(z))\,d\mu^{|Z^{n}}(z)\right].$
We end with a generalization of Theorem 2 for the fuzzy setting. We omit its proof, as it is analogous to that of Theorem 2.
Theorem 5.
Let $\mu^{|Z^{n}}$ be a $Z^{n}$-dependent probability kernel, and $U:[0,\infty]\to[-\infty,\infty]$. Then $\mathcal{E}^{*}$ maximizes
$\mathrm{E}^{P_{*}^{Z^{n}}}\left[\int_{\mathcal{Z}}U(\mathcal{E}^{Z^{n}}(z))\,d\mu^{|Z^{n}}(z)\right],$
if and only if $\mathcal{E}$ is the $U$-expected-utility optimal e-value against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
Remark 5 (Choice of utility function).
In general, the utility $U(e)$ of a realization $e$ of our e-value expresses how much we value a rejection at level $e=1/\alpha$. In the e-value literature, the most popular utility is log-utility $U:x\mapsto\log(x)$ (grunwald2023safe; larsson2024numeraire). The log-utility function may be interpreted as valuing evidence linearly in the order of magnitude: $\log(10)$ is half as valuable as $\log(100)=2\log(10)$. The main motivation for log-utility usually comes from i.i.d. sequential settings, where log-utility optimality can be coupled to minimizing expected stopping times through Wald’s Identity.
In the fuzzy prediction context, such sequential arguments seem less important, so that we see no direct reason to prefer log-utility over other utilities, beyond it serving as a well-studied ‘default’ option. Luckily, more general utilities have also been studied (koning2025continuoustestingunifyingtests). One example are power-utilities $U:x\mapsto(x^{h}-1)/h$, $h\neq 0$, where the log-utility appears as the limit as $h\to 0$, or bounded versions of such utilities.
4.4 Merging fuzzy prediction sets
Fuzzy prediction sets inherit the desirable merging properties of their underlying e-values.
If we have two arbitrary (possibly highly dependent) fuzzy prediction sets $\mathcal{E}_{1}^{Z^{n}}$ and $\mathcal{E}_{2}^{Z^{n}}$ for $Z_{n+1}$, then their average is still a fuzzy prediction set. Indeed, their underlying e-values, say $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, still average to a valid e-value:
$\overline{\mathcal{E}}=(\mathcal{E}_{1}+\mathcal{E}_{2})/2.$
The resulting prediction set $\overline{\mathcal{E}}^{Z^{n}}$ is then a fuzzy prediction set for $Z_{n+1}$. In fact, this is equivalent to simply averaging the prediction sets $\mathcal{E}_{1}^{Z^{n}}$ and $\mathcal{E}_{2}^{Z^{n}}$ for $Z_{n+1}$. This extends from averaging a finite number of prediction sets to mixtures over collections of prediction sets. wang2025onlyadmissible recently proved that (weighted) averaging (possibly with the constant e-value $\mathcal{E}\equiv 1$) is the only admissible way to merge arbitrarily dependent e-values.
Such arbitrarily dependent fuzzy prediction sets may be different fuzzy prediction sets produced on the same data. For example, this enables us to construct an optimal fuzzy prediction set for many choices of $Q$ and produce a ‘mixture’ fuzzy prediction set that weights our preference over options of $Q$. Or it allows us to average over split-conformal fuzzy prediction sets over different choices of the split (vovk2024conformal).
As non-fuzzy prediction sets (possibly of different levels) can be viewed as special fuzzy prediction sets, they can also be merged in the same manner, but the outcome will generally be fuzzy merged prediction sets. Rounding down such a merged fuzzy prediction set to a non-fuzzy prediction set usually discards a lot of evidence.
Another popular method to merge e-values is through multiplication, which is valid under independence. While averaging fuzzy prediction sets cannot increase evidence beyond the maximum evidence for any of the individual prediction sets, multiplying them does have this potential. At the same time, multiplication may also kill-off evidence if one of the fuzzy prediction sets equals zero at some points. Non-fuzzy prediction sets suffer from this issue, as they have zero-values by construction. The assumption of independence may be dramatically weakened to a kind of conditional mean-independence property of one of the fuzzy prediction sets, given the other: $\mathrm{E}[\mathcal{E}_{2}\mid\mathcal{E}_{1}]\leq 1$.
4.5 Example: fuzzy Gaussian prediction
Figure 4: Optimal fuzzy prediction sets for Neyman–Pearson-utility (solid), log-utility (dotted) and bounded log-utility (dashed) under the simple Gaussian setting (left) and composite Gaussian setting (right) from Example 1, 2 and 5, for $n=3$, $\mu=0$, $\sigma=1$, $\tau=3.5$, $\bar{Z}_{n}=1.44$. Here, the Neyman–Pearson utility and bounded log-utility are both bounded at $1/0.05=20$.
Example 5 (Fuzzy Gaussian).
In Section 3.5, we implicitly considered optimal fuzzy prediction sets under the classical Neyman–Pearson utility $x\mapsto x\wedge 1/\alpha$. We continue these examples under more general utility functions.
We first consider the log-utility version of Example 1, where we had the simple model $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$ and optimized against the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for known $\mu\in\mathbb{R}$ and $\tau>\sigma>0$. Here, we could choose $\tau\to\infty$ to mimic the Lebesgue measure, as in Section 3.5, but we find that this generally yields undesirable fuzzy prediction sets.
It is well-known that the log-utility-optimal e-value is the likelihood ratio itself. Hence, the log-utility-optimal e-value here is simply the likelihood ratio
$\mathcal{E}(z^{n+1})=\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})=\mathcal{E}^{z^{n}}(z_{n+1}).$ (12)
For $1/\alpha$-bounded log-utility $U:x\mapsto\log(\mathcal{E}\wedge 1/\alpha)$, the optimal e-value is the likelihood ratio capped at $1/\alpha$ and ‘boosted’ by some constant $b_{\alpha}$ so that it has expectation exactly 1 (koning2025continuoustestingunifyingtests):
$\mathcal{E}^{z^{n}}(z_{n+1})=\left(b_{\alpha}\frac{d\mathcal{N}(\mu,\tau^{2})}{d\mathcal{N}(\mu,\sigma^{2})}(z_{n+1})\right)\wedge 1/\alpha.$
In the left panel of Figure 4, we display these fuzzy prediction sets next to the traditional prediction set from Example 1.
For the composite setting introduced in Example 2, we had derived the likelihood ratio
$\textnormal{LR}(A,B)=\frac{d\mathcal{N}(0,\tau^{2}+\tau^{2}/n)}{d\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)}(B),$
where $A=Z^{n}-\bar{Z}_{n}$ and $B=Z_{n+1}-\bar{Z}_{n}$. Fixing $\bar{Z}_{n}$ and evaluating this likelihood ratio at plug-in values $z_{n+1}\in\mathbb{R}$ for $Z_{n+1}$, we recover the log-optimal fuzzy prediction set:
$\mathcal{E}^{Z^{n}}(z_{n+1})=\frac{d\mathcal{N}(\bar{Z}_{n},\tau^{2}+\tau^{2}/n)}{d\mathcal{N}(\bar{Z}_{n},\sigma^{2}+\sigma^{2}/n)}(z_{n+1}),$ (13)
which relies on a variant of the Hunt-Stein theorem derived for log-optimal e-values proven by perez2022statistics. To the best of our knowledge, this has not been proven yet for general expected-utility-optimal e-values, but we would be surprised if it does not go through in general.
Comparing the fuzzy prediction set in (13) to (12), we see that this likelihood ratio is centered at $\bar{Z}_{n}$ instead of $\mu$ and the variance in both numerator and denominator is inflated by an additional $\tau^{2}/n$- and $\sigma^{2}/n$-term due to the estimation of $\mu$ by $\bar{Z}_{n}$. We illustrate this fuzzy prediction set in the right panel of Figure 4, along its bounded version and the Neyman–Pearson-utility variant from Example 2.
5 Optimal conformal prediction
In this section, we specialize our methodology to conformal prediction, where $\mathcal{P}$ is the class of exchangeable distributions on $\mathcal{Z}^{n+1}$. We cover both optimality for classical conformal prediction sets and more general fuzzy conformal prediction sets.
As we showed in Section 4.2, a valid fuzzy conformal prediction set for the model $\mathcal{P}$ corresponds to a valid e-value $\mathcal{E}$ for ‘hypothesis’ $\mathcal{P}$. Moreover, in Section 4.3 we show that the fuzzy prediction set is optimal if $\mathcal{E}$ is optimal for some expected-utility target $\mathrm{E}^{Q}[U(\mathcal{E})]$. As a consequence, we may reduce the problem of studying optimal (fuzzy) conformal prediction sets to studying optimal e-values for exchangeability. Optimal e-values for exchangeability (and more general group invariance) have recently been developed in koning2023post.
5.1 Background: orbits & exchangeability
Let $\Pi(n+1)$ denote the group of permutations on $n+1$ elements. This group acts on $\mathcal{Z}^{n+1}$ by permuting the elements of each tuple $(z_{1},\dots,z_{n+1})\in\mathcal{Z}^{n+1}$. The group action partitions $\mathcal{Z}^{n+1}$ into a set $\mathcal{O}^{n+1}$ of disjoint orbits $O\in\mathcal{O}^{n+1}$. For a given $z^{n+1}\in\mathcal{Z}^{n+1}$, its orbit is the set of all its permutations:
$O(z^{n+1})=\{z\in\mathcal{Z}^{n+1}:z=\pi z^{n+1},\textnormal{ for some }\pi\in\Pi(n+1)\}.$
On each orbit, we designate a single tuple as its orbit representative, and we define the map $\cdot$] as the map that carries any point on the orbit to this representative.
We say that a random variable $Z^{n+1}$ on $\mathcal{Z}^{n+1}$ is exchangeable if
$Z^{n+1}\overset{d}{=}\pi Z^{n+1},\textnormal{ for every }\pi\in\Pi(n+1).$
Note that this is actually a statement about its distribution $P^{Z^{n+1}}$, which we say is exchangeable if it is the law of an exchangeable random variable. We use $\mathcal{P}^{\textnormal{exch.}}$ to denote the collection of all exchangeable distributions on $\mathcal{Z}^{n+1}$.
Throughout, we use the following well-known result.
Lemma 2.
$P$ is exchangeable if and only if $P^{|O}=\textnormal{Unif}(O)$, for $P^{O}$-almost every orbit $O\in\mathcal{O}^{n+1}$.
5.2 Optimal fuzzy conformal
The key idea in koning2023post to obtain an optimal e-value for exchangeability is to simplify the problem to finding a ‘locally’ optimal e-value $\mathcal{E}_{O}:O\to[0,\infty]$ on each orbit $O\in\mathcal{O}^{n+1}$, and subsequently recombining these ‘local’ e-values into a ‘globally’ optimal e-value:
$\mathcal{E}(Z^{n+1})=\mathcal{E}_{O(Z^{n+1})}(Z^{n+1}).$
Indeed, Theorem 6, which is specialized from Theorem 5 in koning2023post, shows that if each local e-value is optimal, then the resulting unconditional e-value $\mathcal{E}$ is optimal.
Theorem 6 (koning2023post).
If $\mathcal{E}_{O}$ is valid for $\textnormal{Unif}(O)$ and optimal for $\mathrm{E}^{Q^{|O}}[U(\cdot)]$, for every $O$, then $\mathcal{E}$ is valid for $\mathcal{P}^{exch.}$ and optimal against $\mathrm{E}^{Q}[U(\mathcal{E})]$. Moreover, this e-value $\mathcal{E}$ is optimal uniformly in mixtures over $Q^{|O}$.
Theorem 6 reduces the problem to constructing a local e-value $\mathcal{E}_{O}$ for $\textnormal{Unif}(O)$ against $Q^{|O}$. To find such a local e-value, we use the “Neyman–Pearson lemma for e-values” recently derived in koning2025continuoustestingunifyingtests. The classical Neyman–Pearson lemma corresponds to the choice $U(x)=x\wedge 1/\alpha$. The version we present in Theorem 7 specializes to the structure at hand: $O$ is finite and $Q^{|O}$ is absolutely continuous with respect to $\textnormal{Unif}(O)$, since $\textnormal{Unif}(O)$ has full support. Corollary 1 presents a more easily interpretable version under some regularity conditions on $U$ and $Q^{|O}$.
Theorem 7 (koning2025continuoustestingunifyingtests).
Let $U:[0,\infty]\to[-\infty,\infty]$ be concave, non-decreasing and upper-semicontinuous. Then, the optimization problem
$\sup_{\mathcal{E}}E^{Q^{|O}}[U(\mathcal{E})],\textnormal{ s.t. }\mathcal{E}:\mathcal{X}\to[0,\infty],\ E^{\textnormal{Unif}(O)}[\mathcal{E}]\leq 1,$
admits an optimal solution. Moreover, if $\mathcal{E}^{*}$ is an optimizer, then there exists a normalization constant $\lambda_{O}\geq 0$ such that
$\lambda_{O}\bigg/\frac{dQ^{|O}}{d\textnormal{Unif}(O)}$ $\in\partial U(\mathcal{E}^{*}),$
on $\{Q^{|O}>0\}$ and $\mathcal{E}^{*}=0$ on $\{Q^{|O}=0\}$, where $\partial U$ is the superdifferentiable of $U$.
Corollary 1.
If $Q^{|O}$ has full support on $O$ and $U$ is differentiable with strictly decreasing derivative $U^{\prime}$, then an optimizer is
$\mathcal{E}^{*}=(U^{\prime})^{-1}\left(\lambda_{O}\bigg/\frac{dQ^{|O}}{d\textnormal{Unif}(O)}\right),$
where $(U^{\prime})^{-1}$ denotes the functional inverse of $U^{\prime}$.
A consequence of Theorem 7 is that an optimal local e-value $\mathcal{E}_{O}$ only depends on the data through the orbit-conditional likelihood ratio $\textnormal{LR}^{|O}:=dQ^{|O}/d\textnormal{Unif}(O)$. In Proposition 1, we use the structure of $Q=P^{Z^{n}}\otimes\mu^{|Z^{n}}$ to show that this conditional likelihood ratio only depends on $Z^{n+1}$ through $Z_{n+1}$. Its proof may be found in Appendix A.2.
Proposition 1.
If $Z^{n}$ is exchangeable under $Q$, then
$\textnormal{LR}^{|O}(z^{n+1})$ $=\frac{dQ^{|O}}{d\textnormal{Unif}(O)}(z^{n+1})$
$=\frac{Q^{Z_{n+1}|O}(z_{n+1})}{1/(n+1)}=:\textnormal{LR}^{Z_{n+1}|O}(z_{n+1}).$
A consequence of Proposition 1 is that an expected utility optimal local e-value $\mathcal{E}_{O}^{*}$ does not depend on $Z^{n}$. Indeed, we are implicitly considering the following hypotheses on each orbit:
$H_{0}^{O}$ $:Z_{n+1}\sim\textnormal{Unif}[Z^{n+1}]$
$H_{1}^{O}$ $:Z_{n+1}\sim Q^{Z_{n+1}|O}.$
By extension, a globally optimal e-value $\mathcal{E}^{*}$ only depends on $Z^{n}$ through the orbit $O(Z^{n+1})$.
Remark 6 (Conformal with orbit-level ranks).
As optimal e-values only depend on the position of $Z_{n+1}$ in $Z^{n+1}]$, we may equivalently express all the results in terms of the rank of $Z_{n+1}$ among $Z^{n+1}]$.
To define a rank for arbitrary (non-numerical) data, we may define it relative to the orbit representative $Z^{n+1}]$. In particular, we define $\textnormal{Rank}(Z^{n+1})$ as the index at which each element in the tuple $Z^{n+1}$ needs to be placed to obtain the representative $Z^{n+1}]$ of $O$. This specializes to the conventional rank for data that has a natural order, if $Z^{n+1}]$ is sorted in ascending order.
Let $R_{n+1}^{O}$ denote last element of the tuple $\textnormal{Rank}(Z^{n+1})$, which corresponds to the rank of $Z_{n+1}$ among $Z^{n+1}]$. We may then equivalently express the reduced orbit-level hypotheses as
$H_{0}^{O}$ $:R_{n+1}^{O}\sim\textnormal{Unif}\{1,\dots,n\},$
$H_{1}^{O}$ $:R_{n+1}^{O}\sim Q_{R}^{O},$
where $Q_{R}^{O}$ is some arbitrary orbit-dependent distribution on $\{1,\dots,n\}$.
In the literature on conformal prediction it is popular to express everything in terms of unconditional ranks (sometimes confusingly named p-values). Using such unconditional ranks exactly discards the information about the orbit.
5.3 Illustration: i.i.d. and split-conformal
As the conditional likelihood ratio is central in finding the optimal e-value, we first illustrate what such a conditional likelihood ratio may look like in a popular setting, before we proceed with several examples of utility functions.
In Proposition 2, we consider optimality against an i.i.d. alternative of the form $Q=P^{Z^{n}}\times\mu$ where $P^{Z^{n}}=(P^{Z_{1}})^{n}$. For example, such an alternative may arise by plugging-in an i.i.d. estimator $\widehat{P}^{Z^{n}}=(\widehat{P}^{Z_{1}})^{n}$ for $P_{*}^{Z^{n}}$, based on a separate sample akin to the popular ‘split-conformal prediction’ approach. Its proof is found in Appendix A.3.
Proposition 2.
Suppose that $Q$ is so that $Z^{n+1}$ is independent and $Z^{n}$ is i.i.d., with $Q^{Z_{n+1}}$ and $Q^{Z_{1}}$ mutually absolutely continuous. Then,
$\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})$
$=\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{n+1})\Bigg/\left(\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z_{i})\right).$
Proposition 2 may be interpreted as stating that
$\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\propto\frac{d\mu}{d\widehat{P}^{Z_{1}}}(z_{n+1}),$
where the proportionality is up to some orbit-dependent constant. Viewing $\mu$ as some reference measure, an interpretation of this result is that the optimal e-value is decreasing in the marginal $\widehat{P}^{Z_{1}}$-density at $z_{n+1}$. We believe this makes intuitive sense, as values of $z_{n+1}$ that have low density may be interpreted as being ‘outlying’ or ‘non-conforming’.
5.4 Generalized-mean utilities
In the e-value literature, the near-universal choice of utility is the log-utility $U:x\mapsto\log(x)$, for which the optimal e-value is the orbit-conditional likelihood ratio itself:
$\mathcal{E}^{\log}(z^{n+1})=\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}).$
This generalizes to the power-utility $U:x\mapsto(x^{h}-1)/h$, $h<1$, $h\neq 0$, where the optimizer equals
$\mathcal{E}^{h}(Y_{n+1})=\frac{(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}{\frac{1}{n+1}\sum_{j=1}^{n+1}(\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1}))^{\frac{1}{1-h}}}.$
The log-utility appears as the $h\to 0$ limit.
The parameter $h$ can be interpreted as tuning how ‘risky’ the e-value is, where large values of $h$ (close to 1) yield a risky e-value and small values ($h\to$-$\infty$) a less risky e-value. Indeed, assuming $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}>0$, the $h\to$-$\infty$ limit is the no-risk e-value
$\lim_{h\to-\infty}\mathcal{E}^{h}\equiv 1.$
Moreover, assuming $\textnormal{LR}^{Z_{n+1}|O}$ is uniquely maximized at $z_{*}^{O}$, the $h\to 1$ limit is the most risky e-value that is large on a single element on the orbit and zero elsewhere:
$\lim_{h\to 1}\mathcal{E}^{h}(z^{n+1})=\begin{cases}n+1,&\textnormal{ if }z_{n+1}=z_{*}^{O(z^{n+1})},\\ 0,&\textnormal{ otherwise}.\end{cases}$
5.5 Optimal classical conformal
In Theorem 8, we show how our optimality framework nests classical conformal prediction as a special case if we choose the Neyman–Pearson utility function $U_{\alpha}^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, $\alpha\in(0,1)$. This nests classical conformal prediction for any non-conformity score that is a monotone function of $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}=(n+1)\times Q^{Z_{n+1}|O(z^{n+1})}$.
Theorem 8.
Write $O^{\prime}$ for $O(z^{n+1})$. A $U_{\alpha}^{\textnormal{NP}}$-optimal e-value is
$\mathcal{E}^{\textnormal{NP}}(z^{n+1})=\begin{cases}1/\alpha,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})>c_{\alpha}^{O^{\prime}},\\ k^{O^{\prime}},\hskip-11.38092pt&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})=c_{\alpha}^{O^{\prime}},\\ 0,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})<c_{\alpha}^{O^{\prime}},\end{cases}$
where $c_{\alpha}^{O^{\prime}}$ is the $\alpha$ upper-quantile of $\textnormal{LR}^{|O^{\prime}}(z^{n+1})$ under $\textnormal{Unif}[Z^{n+1}]$, and $k^{O^{\prime}}$ is some orbit-dependent constant that ensures $\mathcal{E}^{\textnormal{NP}}(Z^{n+1})$ is an exact e-value. This e-value is optimal uniformly in any mixture over orbits, and in any monotone transformation of the likelihood ratio.
To make this more interpretable, we provide Corollary 2, which combines Theorem 8 with our result for the i.i.d. setting in Proposition 2. Here, we consider an i.i.d. estimator $\widehat{P}^{Z^{n}}=(\widehat{P}^{Z_{1}})^{n}$ for $P_{*}^{Z^{n}}$, and minimize the expected measure of the prediction set
$\mathrm{E}^{\widehat{P}^{Z^{n}}}[\mu(C_{\alpha}^{Z^{n}})].$
Using $f=d\widehat{P}^{Z_{1}}/d\mu$ to denote the density, Corollary 2 shows that classical conformal prediction is optimal in this sense when the conformity score $s$ is chosen as a monotone function of the density $f$. Conversely, given a conformity score $s$ and measure $\mu$, classical conformal prediction is optimal uniformly in the class of true distributions for which $s$ is a monotone function of the corresponding density $f$.
Corollary 2.
Suppose $\widehat{P}^{Z^{n}}=(\widehat{P}^{Z_{1}})^{n}$, and let $s:\mathcal{Z}\to\mathbb{R}_{+}$ denote a conformity score. Then, classical conformal prediction minimizes the expected prediction set size for any $\mu$ and $\widehat{P}^{Z_{1}}$ that satisfy
$s=m\left(\frac{d\widehat{P}^{Z_{1}}}{d\mu}\right),$
where $m:[0,\infty]\to[-\infty,\infty]$ is some strictly monotone function.
5.6 Clipping, capping and dampening
In this section, we cover some practical tools for designing utility functions.
For subsequent decision making, as treated in Section 6, it may be desirable to place a lower bound $0<b\leq 1$ on our e-values, such as $b=0.01$ or $b=0.1$. This prevents settings in which a single zero-valued e-value may come to dominate subsequent decision making due to the minimax nature of the decisions. The log-optimal e-value with such a lower bound is
$\mathcal{E}(z^{n+1})=\left(\lambda\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\right)\vee b,$ (14)
where $\lambda$ is some normalization constant that ensures $\mathcal{E}$ is an exact e-value. This may be interpreted as ‘clipping’ the likelihood ratio from below, and simultaneously shrinking it by $\lambda$ to ensure it remains a valid e-value. Alternatively, it can be viewed as choosing a utility function with $U(x)=$-$\infty$ on $[0,b)$.
Analogously, in case we are not interested in evidence beyond some value $c\geq 0$, we may also cap our e-values below some upper bound $c\geq 1$. The resulting optimizer is analogous to (14), with $\vee b$ is replaced by $\wedge c$ (koning2025continuoustestingunifyingtests). Capping may also be incorporated in the utility function, and appears in the Neyman–Pearson utility function $U(x)=x\wedge 1/\alpha$.
An alternative strategy to ‘dampen’ e-values is introduced by grunwald2023posterior. He proposes to take an e-value $\mathcal{E}$ and dampen it to the e-value $b+(1-b)\mathcal{E}$ that may be interpreted as a $b$-mixture of the e-value and the constant 1. If $\mathcal{E}$ is the log-utility maximizer, then we find $b+(1-b)\mathcal{E}$ implicitly maximizes the utility function
$U(\mathcal{E})=\log([\mathcal{E}-b]\vee 0).$
6 Certifying subsequent decisions
In this section, we show how (fuzzy) prediction sets may be used in subsequent decision making.
We consider a decision maker who is faced with selecting a decision $d$ in a decision space $\mathcal{D}$ while facing uncertainty captured by a random variable $Z_{n+1}$ in the outcome space $\mathcal{Z}$. When selecting the decision $d$, they would ideally like to minimize the ‘oracle loss’ $L:\mathcal{D}\times\mathcal{Z}\to\mathbb{R}$ associated with the decision under the true outcome of $Z_{n+1}$: $\min_{d\in\mathcal{D}}L(d,Z_{n+1})$. Unfortunately, the outcome of $Z_{n+1}$ is not yet available at the time of decision making. Instead, the decision maker relies on a (fuzzy) prediction set for $Z_{n+1}$ to aid them in their decision making process.
6.1 Prediction sets and as-if decisions
We begin by presenting the case where the decision maker receives a conventional non-fuzzy prediction set, before extending to fuzzy prediction sets. For these non-fuzzy prediction sets, this was recently developed by andrews2025certified and kiyani2025decision.
Suppose the decision maker is given a valid level $(1-\alpha)$ prediction set $C_{\alpha}^{Z^{n}}\subseteq\mathcal{Z}$ for $Z_{n+1}$ under $\mathcal{P}$. We define the as-if decision $\delta(C_{\alpha}^{Z^{n}})$, that pretends as if $Z_{n+1}\in C_{\alpha}$. That is, it minimizes the worst-case loss over all possible realizations of $Z_{n+1}$ in the prediction set $C_{\alpha}^{Z^{n}}$:
$\delta(C_{\alpha}^{Z^{n}}):=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in C_{\alpha}^{Z^{n}}}L(d,z),$ (15)
where we assume the minimum exists (manski2021econometrics). Note that if the prediction set happens to only contain the true future realization of $Z_{n+1}$, then this action would minimize the oracle loss.
kiyani2025decision and andrews2025certified show that this as-if decision $\delta(C_{\alpha}^{Z^{n}})$ enjoys a $(1-\alpha)$-confidence lower bound on the to-be-realized loss:
$P[L(\delta(C_{\alpha}^{Z^{n}}),Z_{n+1})\leq R(\delta(C_{\alpha}^{Z^{n}}))]\geq 1-\alpha,$
for every $P\in\mathcal{P}$, where
$R(C_{\alpha}^{Z^{n}}):=\sup_{z\in C_{\alpha}^{Z^{n}}}L(\delta(C_{\alpha}^{Z^{n}}),z).$
is the worst-case loss associated with this action if $z\in C_{\alpha}^{Z^{n}}$.
6.2 Fuzzy prediction sets
We now move to fuzzy prediction sets, and showcase two distinct ways in which they generalize the as-if decision making framework: as-if decisions under the post-hoc selection of the confidence level $\alpha$ and minimax decisions over a loss function weighted by the fuzzy prediction set.
Minimax decisions over an evidence-weighted loss function have previously been for fuzzy confidence sets on parameter spaces under the name ‘E-posterior’ and ‘E-certification’ (grunwald2023posterior; andrews2025certified). While post-hoc level selection has been explored in the context of conformal prediction before (gauthier2025values), the idea of post-hoc as-if decisions is novel, to the best of our knowledge. Moreover, we are also the first to consider such weighted decisions outside of a parameter space.
6.2.1 Post-hoc level as-if decisions
A downside of subsequent decisions with non-fuzzy prediction sets, as presented in Section 6.1, is that the decision maker must pre-specify a single confidence level $\alpha$. In this section, we show that a fuzzy prediction set $\mathcal{E}^{Z^{n}}$ enables a decision maker to choose among as-if decisions at each confidence level $\alpha$, and select the decision with the desired confidence level $\alpha$ in an arbitrary data-dependent fashion. Overall, this still provides a post-hoc $\alpha$ confidence guarantee.
In particular, we first construct the sublevel sets
$\overline{C}_{\alpha}^{Z^{n}}$ $=\left\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\right\}.$
For each sublevel set $\overline{C}_{\alpha}$, we then construct the as-if decision
$\delta(\overline{C}_{\alpha}^{Z^{n}})=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in\overline{C}_{\alpha}^{Z^{n}}}L(d,z).$
This produces a collection $(\delta(\overline{C}_{\alpha}^{Z^{n}}))_{\alpha>0}$ of possible decisions. A decision maker may then select the desired level of confidence $\widetilde{\alpha}$ in an arbitrary data-dependent fashion. The resulting associated decision then satisfies the confidence guarantee
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P\left[L(\delta(\overline{C}_{\widetilde{\alpha}}^{Z^{n}}),z)\geq R(\delta(\overline{C}_{\widetilde{\alpha}}^{Z^{n}}))\mid\widetilde{\alpha}\right]}{\widetilde{\alpha}}\right]\leq 1,$ (16)
for every data-dependent $\widetilde{\alpha}$ and $P\in\mathcal{P}$, mimicking the confidence guarantee treated in Theorem 3.
As a traditional (non-fuzzy) level $\alpha$ prediction set may be seen as a special binary fuzzy prediction set, we may technically also use data-dependent levels with traditional prediction sets. However, we then find that if we select $\widetilde{\alpha}>\alpha$ this yields the same prediction set as $\alpha$ but with a weaker guarantee $\widetilde{\alpha}$, and if $\widetilde{\alpha}<\alpha$ then the resulting prediction set is the entire sample space. In this sense, a non-fuzzy level $\alpha$ prediction set effectively forces the choice $\widetilde{\alpha}=\alpha$ as the only sensible option.
6.2.2 Weighted loss
Instead of producing a collection of possible as-if decisions under a post-hoc level guarantee, it is also possible to weight our loss by the values of the fuzzy prediction set.
In particular, we may consider the decision
$\delta(\mathcal{E}^{Z^{n}}):=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in\mathcal{Z}}L(d,z)/\mathcal{E}^{Z^{n}}(z),$
where we downweight values of $z$ against which we have much evidence, as such values are unlikely to be the realization of $Z_{n+1}$. This puts most emphasis on values of $z$ that may plausibly be the realization of $Z_{n+1}$.
The corresponding risk bound is
$R(\mathcal{E}^{Z^{n}}):=\sup_{z\in\mathcal{Z}}L(\delta(\mathcal{E}^{Z^{n}}),z)/\mathcal{E}^{Z^{n}}(z),$
on which our decision enjoys the confidence bound
$\mathrm{E}^{P}\left[\frac{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})}{R(\mathcal{E}^{Z^{n}})}\right]$
$=\mathrm{E}^{P}\left[\frac{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})}{\sup_{z\in\mathcal{Z}}L(\delta(\mathcal{E}^{Z^{n}}),z)/\mathcal{E}^{Z^{n}}(z)}\right]$
$\leq\mathrm{E}^{P}\left[\frac{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})}{L(\delta(\mathcal{E}^{Z^{n}}),Z_{n+1})/\mathcal{E}^{Z^{n}}(Z_{n+1})}\right]$
$=\mathrm{E}^{P}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1.$
6.3 Connecting as-if and weighted
While the weighted loss approach for fuzzy prediction sets in Section 6.2.2 and the as-if decision approach from Section 6.1 may seem distinct, we show that as-if approach may be viewed as a special limiting case of the weighted loss approach. This relationship was not yet established before; andrews2025certified describe both approaches as distinct paradigms. The connection is not obvious, as it relies on our observation that the E-posterior coined by grunwald2023posterior is equivalent to a fuzzy confidence set, and so a generalization of a confidence set.
Let $\gamma\in(0,1)$, and define the following fuzzy prediction set for every $z_{n+1}\in\mathcal{Z}$:
$\mathcal{E}^{Z^{n}}(z_{n+1}):=\gamma\frac{1}{\alpha}\mathbb{I}\{z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\mathbb{I}\{z_{n+1}\in\mathcal{C}_{\alpha}^{Z^{n}}\}.$ (17)
Note that $\mathcal{E}$ converges to a non-fuzzy prediction set when $\gamma\to 1$. Moreover, this is indeed an e-value because it is dominated by the mixture $\gamma\frac{1}{\alpha}\mathbb{I}\{Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)$ of a valid e-value $\frac{1}{\alpha}\mathbb{I}\{Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}}\}$ and the constant 1, and a mixture of valid e-values is itself a valid e-value.
Plugging this into definition of the risk bound for a given decision $d$ yields
$\sup_{z}\frac{L(d,z)}{\mathcal{E}^{Z^{n}}(z)}=\sup_{z}\frac{L(d,z)}{\gamma\frac{1}{\alpha}\mathbb{I}\{z\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}}$
$=\sup_{z}\left\{\frac{\alpha}{\gamma}\mathbb{I}\{z\not\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)+\frac{1}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)\right\}.$
Now, assuming $\mathcal{C}_{\alpha}^{Z^{n}}$ is non-empty and $L(d,z)$ is bounded, we have for $\gamma$ sufficiently close to 1 that this equals
$\sup_{z}\frac{1}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)=\frac{1}{1-\gamma}\sup_{z\in\mathcal{C}_{\alpha}^{Z^{n}}}L(d,z),$
which is the risk function we minimize for non-fuzzy prediction sets in (15), up to the constant $1/(1-\gamma)$ that is irrelevant for the optimization. This shows that the non-fuzzy ‘as-if’ decision can indeed be viewed as a limiting case of the fuzzy procedure.
7 Covariates
In many applications, the observations $Z^{n+1}$ are decomposed into $Z^{n+1}=(X^{n+1},Y^{n+1})$, where the outcomes $Y^{n+1}$ come with covariates $X^{n+1}$. In such a setting, the covariate $X_{n+1}$ is observed alongside $Z^{n}=(X^{n},Y^{n})$, and we are to construct a prediction set for $Y_{n+1}$. In this section, we show how such covariates are easily incorporated into our framework. We only cover the non-fuzzy setting, as fuzzy prediction sets may be derived analogously.
When adding covariates, our prediction set $C_{\alpha}^{(Z^{n},X_{n+1})}$ now also depends on $X_{n+1}$. The coverage guarantee is also marginal over the covariates:
$P(Y_{n+1}\in C_{\alpha}^{(Z^{n},X_{n+1})})\geq 1-\alpha,$
for every $P\in\mathcal{P}$.
Following the discussion in Section 3.1, we may still construct a prediction set $C_{\alpha}$ for $Z^{n+1}$, and subsequently slice it at the realization $(Z^{n},X_{n+1})=(z^{n},x_{n+1})$ to obtain our prediction set $C_{\alpha}^{(z^{n},x_{n+1})}$ for $Y_{n+1}$:
$C_{\alpha}^{(z^{n},x_{n+1})}:=\{y\in\mathcal{Y}:C_{\alpha}(z^{n},(x_{n+1},y))=0\}.$
The optimality results remain the same, but we now maximize the power under the distribution
$Q=P_{*}^{Z^{n},X_{n+1}}\otimes\mu^{|(Z^{n},X_{n+1})}.$
7.1 Conformal prediction with covariates
In the application to the conformal setting, the only thing that changes is that we may use some additional structure in $Q$. In particular, we must replace $Q^{Z_{n+1}}=\mu^{|Z^{n}}$ by $Q^{Z_{n+1}}=P_{*}^{X_{n+1}|Z^{n}}\otimes\mu^{|(X_{n+1},Z^{n})}$
For example, the i.i.d. result in Proposition 2 now specializes to
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})$ $\propto\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z^{n+1})$
$=\frac{d(P_{*}^{X_{n+1}}\otimes\mu^{|X_{n+1}})}{dP_{*}^{X_{n+1},Y_{n+1}}}(z_{n+1})$
$=\frac{dP_{*}^{X_{n+1}}\otimes d\mu^{|X_{n+1}}}{dP_{*}^{X_{n+1}}\otimes dP_{*}^{Y_{n+1}|X_{n+1}}}(z_{n+1})$
$=\frac{d\mu^{|X_{n+1}}}{dP_{*}^{Y_{n+1}|X_{n+1}}}(y_{n+1}),$
where the normalization constant is
$\frac{1}{(n+1)}\sum_{i=1}^{n+1}\frac{d\mu^{|X_{i}}}{dP_{*}^{Y_{n+1}|X_{i}}}(y_{i}).$
Here, we may write $P_{*}^{Y_{n+1}|X_{i}}$ instead of $P_{*}^{Y_{i}|X_{i}}$ due to the i.i.d. assumption.
To apply this, we may plug-in some estimator $\widehat{P}^{Y_{n+1}|X_{n+1}}$, for the kernel $P_{*}^{Y_{n+1}|X_{n+1}}$, or several estimators and average the resulting e-values. We use this in Section 8, where we estimate this kernel through a neural network.
8 Application: image recognition
In this section, we showcase our methodology in a classical Machine Learning problem: character recognition. Here, we use the same data and model as considered by gauthier2025values.
In particular, we consider the Federated Extended MNIST (FEMNIST) dataset (caldas2018leaf). In this dataset, $Y_{i}$ is one of 62 possible characters (a-z, A-Z, 0-9) and $X_{i}$ is a $28\times 28$ pixel image of a handwritten version of this character. We split the data into a training set (80%), calibration set (15%) and test set (5%), where the percentages are approximate because we put characters by the same writer into the same subset.
We use the training data to ‘train’ the same LeNet-inspired (lecun1998mnist) neural network as used by gauthier2025values, which is an estimator of the kernel $P_{*}^{Y_{n+1}|X_{n+1}}$. We plug this estimator into the likelihood ratio described in Section 7.1, and subsequently use this to construct several utility-optimal e-variables, as described in Section 5. The utilities we consider are all special cases of the capped power-utility $U:x\mapsto((x\wedge 1/\alpha)^{h}-1)/h$ framework introduced by koning2025continuoustestingunifyingtests, which interpolates between log-optimal e-value ($\alpha=0$ and $h\to 0$) and Neyman–Pearson-optimal e-value ($\alpha>0$ and $h=1$). We compute the fuzzy prediction set by taking the calibration set to be $Z^{n}$, and we select the image of a single observation from the test set to be $X_{n+1}$, so that our fuzzy prediction sets are on the true label $Y_{n+1}$.
In Figure 5, we plot the resulting fuzzy prediction sets for the label $Y_{n+1}$ of the image $X_{n+1}$ given in Figure 6. Here, each bar represents the amount of evidence (e-value) against every character. The plots are sorted from least-to-most evidence, on which they all agree (up to ties) by construction. Here, we see that the parameters $\alpha$ and $h$ are jointly able to change the shape of the fuzzy prediction set quite dramatically.
The approximate staircase-shape of the capped power-utility plot (bottom-left) makes it a good starting point to explain how we may interpret these plots. Ignoring the slightly elevated evidence at characters $G$, $b$ and $a$, this fuzzy prediction set can be interpreted as simultaneously reporting multiple prediction sets for different data-dependent level certificates $\widetilde{\alpha}$:
[…]
We may contrast this to the Neyman–Pearson plot (bottom-right), for which this becomes:
$C_{\widetilde{\alpha}}=\begin{cases}\{d,0,G,b\},&\textnormal{ for }0.001\leq\widetilde{\alpha},\\ \textnormal{all characters},&\textnormal{ for }\widetilde{\alpha}<0.001.\end{cases}$
This shows that a non-fuzzy prediction set yields a much more refined and non-binary expression of the available evidence. Comparing the two, a benefit of this fuzzy prediction set is that it reports a prediction set $\{d\}$ at a certificate of level $0.01$, which is unavailable for the Neyman–Pearson prediction set. The cost is that $\{d,0,G,b\}$ only comes with a certificate of $0.00125$, instead of the $0.001$ offered by the Neyman–Pearson variant.
The capped log-utility (top-right) and especially the uncapped log-utility (top-left) show a much wider spectrum of evidence. The uncapped log-utility displays strong evidence against characters such as $9$ and $R$, which is not visible in the other fuzzy prediction sets that are capped at $1000$. This may be of interest if the subsequent task is to eliminate certain characters from contention, instead of determining the correct character. Indeed, the log-utility maximizer reports at least 13 characters for post-hoc significance levels $\widetilde{\alpha}<0.1$, which does not seem helpful if we desire to identify the ‘correct’ character. This suggests that log-utility, which is often pushed as ‘the right utility function’ in the e-value literature, is not necessarily ideal for constructing fuzzy prediction sets. At the same time, such behavior may be attractive in different applications, such as a preliminary medical diagnosis to help rule out several implausible diseases.
An important open question is what kind of utility functions give rise to desirable fuzzy prediction sets for certain subsequent decision tasks. For now, the parameters $\alpha$ and $h$ seem like two useful instruments to shape our utility function, and thereby the resulting prediction sets: $\alpha$ caps the maximum evidence that we are interested in, and $h\in[-\infty,1]$ influences how ‘risky’ or ‘aggressive’ the e-variable is.
Figure 5: Expected-power-utility-optimal $0,1/\alpha]-$valued fuzzy prediction sets over possible labels, for $h=0$ (log-utility), $h=1/2$ (power-utility) and $h=1$ (Neyman–Pearson), for values $\alpha=0$ (top-left) and $\alpha=0.001$ (others). The true character label is marked in red (d).
Figure 6: Handwritten character used in the application.
9 Data availability
Code to replicate the application may be found at https://github.com/nickwkoning/fuzzy_prediction_sets.
10 Acknowledgements
We thank Etienne Gauthier, Peter Grünwald, Guneet Dhillon, Yash Nair and Junu Lee for their feedback and discussions.
Appendix A Omitted proofs
A.1 Proof of Theorem 3
Proof.
For each $\overline{C}_{\alpha}$, (8) reveals that the smallest data-dependent level $\widetilde{\alpha}$ choice for which $Z_{n+1}$ is excluded is $1/\mathcal{E}^{Z^{n}}(Z_{n+1})$. This implies,
$\frac{\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\leq\frac{\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}\}}{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}=\mathcal{E}^{Z^{n}}(Z_{n+1}),$
for every $\widetilde{\alpha}$. As a consequence,
$\mathrm{E}_{\widetilde{\alpha}}^{P}$ $\left[\frac{P(Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right$]
[…]
$=\mathrm{E}^{P}\left[\frac{\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\right]\leq\mathrm{E}^{P}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1.$
A.2 Proof of Proposition 1
To prove Proposition 1, we first prove the following lemma.
Lemma 3.
If $Z^{n}$ is exchangeable under $Q$ then
$Q^{Z^{n}|O,Z_{n+1}}=\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}),$ (18)
where with $\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\})$ we mean uniformly sampled without replacement from the tuple $Z^{n+1}]$ with $Z_{n+1}$ removed.
Proof.
If $Z^{n+1}$ is exchangeable, then $Z^{n}$ is exchangeable. Exchangeability of $Z^{n}$ means that the distribution of $Z^{n+1}$ is invariant under permutations in $\Pi(n)$, which is equivalent to uniformity on each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$, as $\Pi(n)$ is a compact group. Each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$ is simply the subset of a $\Pi(n+1)$-orbit $O\in\mathcal{O}^{n+1}$ in which the $(n+1)$th observation is fixed. In fact, each $\Pi(n+1)$-orbit may be partitioned into $\Pi(n)$-orbits as subsets, each differing by the $(n+1)$th element. Hence, conditionally on $O\in\mathcal{O}^{n+1}$, each conditional distribution $Q^{|O}$ may be viewed as first sampling this $(n+1)$th element using some distributions (and with it the $\Pi(n)$-orbit), and subsequently sampling uniformly from the $\Pi(n)$-orbit that contains this element. ∎
Proof of Proposition 1.
We have
$\textnormal{LR}^{|O}(z^{n+1})=\frac{dQ^{|O}}{d\textnormal{Unif}(O)}(z^{n+1})$
$=\frac{d(Q^{Z_{n+1}|O}\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}{d(\textnormal{Unif}[Z^{n+1}]\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}(z^{n+1})$
$=\frac{dQ^{Z_{n+1}|O}}{d\textnormal{Unif}[Z^{n+1}]}(z_{n+1})=(n+1)\times Q^{Z_{n+1}|O}(z_{n+1})$
$=:\textnormal{LR}^{Z_{n+1}|O}(z_{n+1}).$
A.3 Proof of Proposition 2
Proof.
By assumption, $Q=Q_{1}\times Q_{1}\times\cdots\times Q_{1}\times Q_{n+1}$. Define $Q^{(i)}$ be this product measure, but with $Q_{n+1}$ in the $i$th position, and $Q_{1}$ in the remaining positions. By the structure of $Q$,
$\overline{Q}=\frac{1}{|\Pi(n+1)|}\sum_{\pi\in\Pi(n+1)}\pi Q=\frac{1}{n+1}\sum_{i=1}^{n+1}Q^{(i)}.$
By Proposition 3 in koning2023post, the conditional likelihood ratio is the restriction of
$\frac{dQ}{d\overline{Q}}$
to $O$. By definition of $\overline{Q}$, we have, for every event $A$,
$\overline{Q}(A)$ $=\frac{1}{n+1}\sum_{i=1}^{n+1}Q^{(i)}(A)$
$=\frac{1}{n+1}\sum_{i=1}^{n+1}\int_{A}\frac{dQ^{(i)}}{dQ}\,dQ$
$=\int_{A}\left(\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{(i)}}{dQ}\right)\,dQ,$
so that
$\frac{d\overline{Q}}{dQ}=\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{(i)}}{dQ}.$
Taking the reciprocal gives
$\frac{dQ}{d\overline{Q}}=\frac{1}{\frac{1}{n+1}\sum_{i=1}^{n+1}\frac{dQ^{(i)}}{dQ}}.$ (19)
Next, by definition of $Q$ and $Q^{(i)}$, we have
$\frac{dQ^{(i)}}{dQ}(z^{n+1})$ $=\frac{\tfrac{dQ_{n+1}}{dQ_{1}}(z_{i})}{\tfrac{dQ_{n+1}}{dQ_{1}}(z_{n+1})}$
Substituting this into (19) and rewriting yields
$\frac{dQ}{d\overline{Q}}(z^{n+1})=\frac{\tfrac{dQ_{n+1}}{dQ_{1}}(z_{n+1})}{\frac{1}{n+1}\sum_{i=1}^{n+1}\tfrac{dQ_{n+1}}{dQ_{1}}(z_{i})}.$